### v1

# Error Correction Code Transformer: From Non-Unified to Unified

Abstract
Channel coding is vital for reliable data transmission in modern wireless systems, and its significance will increase with the emergence of sixth-generation (6G) networks, which will need to support various error correction codes. However, traditional decoders were typically designed as fixed hardware circuits tailored to specific decoding algorithms, leading to inefficiencies and limited flexibility. To address these challenges, this paper proposes a unified, code-agnostic Transformer-based decoding architecture capable of handling multiple linear block codes, including Polar, Low-Density Parity-Check (LDPC), and Bose-Chaudhuri-Hocquenghem (BCH), within a single framework. To achieve this, standardized units are employed to harmonize parameters across different code types, while the redesigned unified attention module compresses the structural information of various codewords. Additionally, a sparse mask, derived from the sparsity of the parity-check matrix, is introduced to enhance the model’s ability to capture inherent constraints between information and parity-check bits, resulting in improved decoding accuracy and robustness. Extensive experimental results demonstrate that the proposed unified Transformer-based decoder not only outperforms existing methods but also provides a flexible, efficient, and high-performance solution for next-generation wireless communication systems.
1 Introduction
With the continuous evolution of wireless communication technologies, the imminent arrival of sixth-generation (6G) wireless communications heralds a paradigm shift towards ultra-reliable, low-latency communications and notably enhanced data rates [1, 2, 3]. To fulfill the ambitious performance requirements of 6G networks, probabilistic coding schemes are anticipated to play a crucial role. In particular, linear block codes such as Polar codes [4] and Low-Density Parity-Check (LDPC) codes [5] are expected to maintain their competitive edge established in fifth-generation (5G) wireless communications. These codes have also emerged as promising candidates for efficient error correction in 6G [6]. Additionally, another algebraic linear block code, known as the Bose-Chaudhuri-Hocquenghem (BCH) code [7], is often used as an outer code to enhance decoding performance.
1.1 Prior Works
Looking at 5G and earlier wireless communication systems, linear block code decoders were typically designed as fixed hardware circuits tailored to specific decoding algorithms. Looking ahead to 6G, this trend is likely to persist [8]. For instance, the Successive Cancellation (SC) decoding algorithm [4] for Polar codes uses a sequential method that traverses a decision tree to obtain the decoding result. This approach simplifies the hardware implementation of Polar codes by pruning binary trees [9]. In contrast, the Sum-Product (SP) decoding algorithm [10] for LDPC codes uses iterative message passing, allowing for parallel data processing. This fundamentally differs from the sequential method used for Polar codes in hardware implementation [11]. Furthermore, BCH codes typically use algebraic decoding techniques [12], classified as hard decoding, which differs greatly from soft decoding methods like SC and SP. Overall, these algorithms provide strong error correction capabilities. However, due to the non-unified nature of the sequential and parallel architectures of Polar and LDPC codes, and the gap between hard and soft decoding in BCH codes, unique hardware circuits must be customized and fine-tuned for each specific linear block code. This may lead to significant engineering effort and inefficient use of hardware resources.
In an effort to reduce hardware resource consumption and establish a quasi-unified decoding framework, three methods have been explored: resource sharing, noise estimation, and Artificial Intelligence (AI)-driven approaches. Resource sharing involves time-sharing storage and computing units. Noise estimation methods decode various linear block codes by estimating the noise of received signals. Meanwhile, AI-driven methods offer a convenient way to unfold traditional algorithms like Belief Propagation (BP).
First, resource-sharing methods, such as those in [13] and [14], endeavor to integrate decoders for Polar and LDPC codes into a shared architecture based on traditional decoding algorithms. Both approaches aim to conserve hardware resources by temporally multiplexing the computational and storage units shared by the decoder through reconfiguration. However, due to the distinct decoding algorithms for each code type, achieving a truly unified architecture with such decoders is generally challenging.
Second, noise estimation methods were first proposed in [15] as the Guessing Random Additive Noise Decoding (GRAND) algorithm, used for Maximum Likelihood (ML) decoding of arbitrary codebooks in discrete channels, with or without memory [16, 17]. The algorithm operates by rank-ordering noise sequences based on their likelihood and subtracting them from the received signal to identify codebook members [16]. GRAND is shown to be capacity-achieving with random codebooks. However, this work faces the challenge of excessive guessing required by the GRAND algorithm, particularly for high logistic weights, which can lead to computational inefficiency. For instance, when the logistic weight equals $128$, up to $5.33\times 10^{7}$ queries may be necessary to complete a single decoding process [17].
Third, in recent years, AI-driven decoders have gained attention as an alternative. While model-based AI methods leverage traditional decoder algorithms and rely on neural networks to simulate belief propagation, their performance improvements are often limited by the constraints of conventional models [18], [19]. In contrast, model-free approaches offer a more flexible and scalable solution. Unlike model-based methods, which depend on predefined mathematical models, model-free approaches do not assume prior knowledge of the encoder or decoder. Instead, they rely on large amounts of labeled data to learn the decoding process directly, bypassing the limitations of conventional algorithms [20]. For example, inspired by the success of Transformers in Natural Language Processing (NLP) [21], a Transformer-based, model-free neural decoder called the Error Correction Code Transformer (ECCT) is proposed [22]. ECCT leverages the Transformer’s self-attention mechanism, where multiple attention heads work in parallel to effectively manage sequential data and capture long-range dependencies, to establish relationships between the input Log-Likelihood Ratios (LLRs) in the decoder. The inherent connection between Transformers and decoders lies in their shared ability to translate sequences—whether from one language to another or from LLRs back to the original information bits. However, despite offering greater flexibility, model-free approaches like ECCT still require frequent tuning of model parameters for each code type, length, and rate. This increases training costs and hardware overhead, making the decoding architecture non-unified.
1.2 Contributions
Thus, to reduce engineering efforts, training costs, and hardware consumption, we propose a unified, code-agnostic Transformer-based decoding architecture capable of handling multiple linear block codes, including Polar, LDPC, and BCH, within a single framework. To achieve this, we designed new standardized units and a unified attention module. Additionally, a sparse masking mechanism was developed to enhance decoding performance and stability  [^1] Simulation codes will be provided to reproduce the results in this paper: http://oa.ee.tsinghua.edu.cn/dailinglong/publications/publications.html. Specifically, the main contributions of this paper are summarized as follows.
  1. 1.
    Unified Decoding Architecture: We introduce a unified, code-agnostic Transformer decoding architecture, capable of seamlessly integrating various linear block codes. This architecture features standardized units to align various code lengths and, for the first time, designs a unified attention module that learns to compress the structural information of all codewords, thereby achieving a code-agnostic decoder. This architecture obviates the necessity for distinct hardware circuits for different decoding algorithms, reducing engineering efforts and hardware resource inefficiency.
  2. 2.
    Standardized Units: Standardized units are used to align parameters such as code length and code rate across different code types. Specifically, after receiving the channel output, we pad different codewords and their corresponding syndromes with zeros up to the maximum length. Through this unit, the Transformer neural decoder can accommodate any codeword within the maximum length, allowing for mixed training in a unified manner, thereby reducing training overhead.
  3. 3.
    Unified Attention: We redesigned the self-attention module to create a unified attention mechanism that allows different attention heads to share attention scores. This approach contrasts with traditional self-attention mechanisms, where each head independently computes its scores using separate learned weights. The unified attention compresses the structural information of different codewords through learning and shares parameters across various heads in the Transformer, resulting in a more cohesive decoding architecture. Additionally, we eliminated the need for iterations through the unified attention mechanism, thereby avoiding the detrimental effects of short cycles in Tanner graphs on LDPC belief propagation decoders. As a result, we have unlocked the full potential of LDPC codes, extending their applicability to short code scenarios.
  4. 4.
    Performance Enhancement: By introducing a sparse mask, which leverages the sparsity of the parity-check matrix, we have significantly enhanced the model’s ability to capture inherent constraints between information and parity-check bits. This has led to a substantial improvement in decoding performance and the achievement of a state-of-the-art design.
  5. 5.
    High Computational Efficiency: The unified attention module in our proposed Transformer-based decoding architecture demonstrates linear computational complexity by eliminating the need to project the input into higher dimensions. This results in a significant reduction in complexity from the traditional $O(N^{2})$ to a more efficient $O(N)$, where $N$ represents the sequence length. Additionally, by employing our sparse mask, we further achieve an $83\%$ reduction in computational complexity. This efficiency gain enables faster processing and lower power consumption, making the architecture suitable for real-time applications in next-generation wireless communication systems.
The rest of this paper is organized as follows. Section II provides background knowledge, including the vanilla Transformers, channel encoders, as well as pre- and post-processing involved to address the issue of overfitting. Section III details our proposed unified code-agnostic Transformer. Section IV presents the results of training and inference. Section V analyzes the impact of the proposed optimization algorithms. Finally, Section VI concludes.
Notations: Lowercase and uppercase bold letters represent vectors and matrices, respectively. GF(2) denotes the binary Galois field. $\mathbb{R}$ represents real numbers. $[\cdot]^{T}$ stands for transpose, while $\mathcal{N}(\mu,{\sigma}^{2})$ signifies a Gaussian distribution with mean $\mu$ and variance ${\sigma}^{2}$. We define $bipolar(\cdot)$ to map binary to bipolar signals, and $bin(\cdot)$ represents its inverse mapping, where $bipolar(x)$ is defined as $1-2x$, and $x$ takes values in the set $\{0,1\}$. $size(\cdot)$ represents the number of elements in a matrix or vector. The sign function is denoted as $sign(\cdot)$.
2 Background
For a better understanding of the proposed unified error correction code Transformer, this section will provide background knowledge on vanilla Transformers and Forward Error Correction (FEC) codes.
2.1 Vanilla Transformers
The vanilla Transformer functions as a sequence-to-sequence model comprising an encoder and a decoder, both built from a series of $L$ identical blocks. Each encoder block primarily integrates a multi-head self-attention module and a position-wise feed-forward network (FFN). To construct a deeper model and prevent gradient degradation, residual connections[23] are employed between the self-attention module and the FFN. Layer normalization[24] is applied to the inputs of each module, enhancing network stability and convergence speed. The overall architecture of the single-layer vanilla Transformer encoder is illustrated in Figure 1.
Figure 1: Overall architecture of the single-layer vanilla Transformer.
The following subsections will introduce the key modules within the vanilla Transformer.
2.1.1 Word Embedding
Word embedding is an important term in natural language processing, used to represent words for text analysis in the form of real-valued vectors. It enhances the computer’s ability to better understand text-based content. In this paper, word embeddings serve as reliability encoders that map the received one-dimensional signal to higher dimensions. It can be expressed as
$Embedding\left(\textbf{{X}}\right)=\textbf{{X}}\odot\textbf{{E}}$ (1)
where $\textbf{{X}}\in\mathbb{R}^{N\times 1}$ is the received signal with length $N$ and dimension one, $\textbf{{E}}\in\mathbb{R}^{N\times d_{k}}$ represents the trainable embedding matrix with dimensions of $d_{k}$, and $\odot$ denotes the Hadamard product.
2.1.2 Multi-Head Self-Attention
The vanilla Transformer establishes an attention mechanism based on the three Query-Key-Value (Q-K-V) matrices. It utilizes the Q and K matrices to establish correlations between different elements within the sequence and weights the V matrix accordingly. The scaled dot-product form of attention is given by
$Attention\left(\textbf{{Q}},\textbf{{K}},\textbf{{V}}\right)=Softmax\left(\frac{\textbf{{QK}}^{T}}{\sqrt{d_{k}}}\right)\textbf{{V}}$ (2)
$Softmax(\textbf{{z}})_{i}=\frac{e^{z_{i}}}{\sum_{j=1}^{n}{e^{z_{j}}}}~{}for~{}i=1,2,\cdots,n$ (3)
where $\textbf{{Q}}\in\mathbb{R}^{N\times d_{k}}$, $\textbf{{K}}\in\mathbb{R}^{M\times d_{k}}$, $\textbf{{V}}\in\mathbb{R}^{M\times d_{v}}$ represent tensors with lengths of $N$ (for queries and keys) and $M$ (for values), and dimensions of $d_{k}$ (for queries and keys) and $d_{v}$ (for values), respectively. The $Softmax\left(\cdot\right)$ function normalizes inputs into the range of 0 to 1, and to alleviate its gradient degradation issue, the dot product of queries and keys is divided by $\sqrt{d_{k}}$.
The study in[21] enhanced the attention layer by introducing a technique known as multi-head attention, which involves $H$ different learned projection sets. This enhancement benefits the attention layer’s performance in two significant ways: it widens the model’s capacity to concentrate on various positions and grants the attention layer multiple representation subspaces. The multi-head attention can be written as
$MultiHeadAttn\left(\textbf{{Q}},\textbf{{K}},\textbf{{V}}\right)=Concat\left(head_{1},head_{2},\cdots,head_{H}\right)\textbf{{W}}^{\textit{O}}$ (4)
where $head_{i}=Attention(\textbf{{QW}}_{i}^{\textit{Q}},\textbf{{KW}}_{i}^{\textit{K}},\textbf{{VW}}_{i}^{\textit{V}})$, and $\textbf{{W}}_{i}^{\textit{Q}}\in\mathbb{R}^{d_{k}\times d_{k}}$, $\textbf{{W}}_{i}^{\textit{K}}\in\mathbb{R}^{d_{k}\times d_{k}}$, $\textbf{{W}}_{i}^{\textit{V}}\in\mathbb{R}^{d_{v}\times d_{v}}$ are trainable parameters, $\textbf{{W}}^{\textit{O}}\in\mathbb{R}^{Hd_{v}\times Hd_{v}}$ is a weight matrix that was trained jointly with the model, serving to condense the outputs of different heads down into a single matrix.
Figure 2: Architecture of multi-head self-attention.
[…]
2.1.3 Position-Wise Feed-Forward Network
The outputs of the self-attention layer are fed into the position-wise FFN layer. The position-wise FFN layer is essentially a fully connected network that operates uniformly on each position, and the non-linear activation function within it enables the Transformer to learn more complex feature representations. The mathematical expression for the FFN is as follows
$FFN\left(\textbf{{X}}\right)=\textbf{{W}}_{2}\times ReLU\left(\textbf{{W}}_{1}\times\textbf{{X}}+\textbf{{b}}_{1}\right)+\textbf{{b}}_{2}$ (5)
$ReLU\left(x\right)=max\left(0,x\right)$ (6)
where X is the output of previous self-attention layer, and $\textbf{{W}}_{1}\in\mathbb{R}^{d_{f}\times d_{h}}$, $\textbf{{W}}_{2}\in\mathbb{R}^{d_{h}\times d_{f}}$, $\textbf{{b}}_{1}\in\mathbb{R}^{d_{f}}$, $\textbf{{b}}_{2}\in\mathbb{R}^{d_{h}}$ are trainable parameters, $d_{h}=H\times d_{k}$ is the dimensions of queries, keys and values, $ReLU\left(\cdot\right)$ represents the activation function. In general, to achieve better performance, $d_{f}$ is set to be greater than $d_{h}$.
2.1.4 Residual Connection and Normalization
The multi-head self-attention layer and the position-wise FFN layer are connected through residual network[23] and layer normalization[24]. The overall architecture of each vanilla Transformer encoder can be written as
$\textbf{{X}}_{1}$ $=LayerNorm\left(MultiHeadAttn(\textbf{{X}})+\textbf{{X}}\right)$ (7)
$\textbf{{X}}_{2}$ $=LayerNorm\left(FFN(\textbf{{X}}_{1})+\textbf{{X}}_{1}\right)$
where $LayerNorm\left(\cdot\right)$ denotes the layer normalization module.
2.2 Forward Error Correction Codes
In information theory, FEC is a technique designed to reduce the bit error rate (BER) when transmitting data over a noisy channel. We establish a standard linear block code utilizing a generator matrix G with dimensions $k\times n$ and a parity-check matrix H with dimensions $(n-k)\times n$. These matrices are specifically designed to satisfy $\textbf{G}\textbf{H}^{T}=\textbf{0}$ over the binary Galois field (GF).
We will now briefly discuss the encoder. The message vector $\textbf{{m}}\in\{0,1\}^{k}$ multiplied with the generator matrix G over the GF(2) results in a codeword vector $\textbf{{x}}\in\{0,1\}^{n}$ satisfying $\textbf{H}\textbf{{x}}=\textbf{0}$. The encoded codeword x is modulated by Binary Phase Shift Keying (BPSK) to $\textbf{{x}}_{s}$, where $BPSK(x)=1-2x$. $\textbf{{x}}_{s}$ is transmitted through binary-input discrete memoryless channels (B-DMCs, e.g. an Additive White Gaussian Noise channel), resulting in a channel output $\textbf{{y}}=\textbf{{x}}_{s}+\textbf{{n}}$, where n is Gaussian noise with zero mean and variance $\sigma^{2}$.
Figure 3: Architecture of the communication system.
The goal of the decoder is to recover the transmitted message m from the received signal y. The overall architecture of the communication system including both encoding and decoding is illustrated in Figure 3. The objective of our work is to design and training of the parameterized decoding function $\textit{f}_{\theta}$.
Figure 4: Decoder framework with pre- and post-processing.
To build a model-free Transformer decoding model, we draw inspiration from the pre- and post-processing methods proposed in [25], which have been demonstrated not to result in performance loss concerning bit error rate or mean squared error (MSE) metrics, as depicted in Figure 4. By training with an all-zero codeword to learn the noise, it is possible to avoid the overfitting problem that arises from the exponential growth of the codeword space with code length. And the B-DMCs is modeled by (8) which is an equivalent statistical mode that differs from the true physical one [26].
$\textbf{{y}}=\textbf{{x}}_{s}\cdot\textbf{{z}}$ (8)
where z is a random multiplicative noise independent of ${\textbf{{x}}_{s}}$.
In the pre-processing step, the absolute values of received signal and the syndromes are concatenated as the input to the neural network, which can be written as
$=\left[\lvert\textbf{{y}}\rvert,s\left(\textbf{{y}}\right)\right]$ y ) ] (9)
[…]
In the post-processing step, the predicted multiplicative noise $\hat{\textbf{{z}}}$ is multiplied by the received signal y to recover x. That is, the predicted $\hat{\textbf{{x}}}$ takes the form
$\hat{\textbf{{x}}}=bin\left(sign\left({\textbf{{y}}}\cdot{\hat{\textbf{{z}}}}\right)\right)$ (10)
3 Proposed Unified Error Correction Code Transformer
In this section, we present the complete architecture and training process of the proposed Unified Error Correction Code Transformer (UECCT).
3.1 Unified Attention
The vanilla self-attention module can be viewed as establishing correlations among different input features within a data sample while neglecting the intrinsic relationships between different samples. This limitation hinders its ability to learn the distinctiveness of features between various code types in channel decoding tasks. Furthermore, research in [27] suggests that the output of the softmax function is probabilistically sparse, indicating that only a few weights dominate. This implies that an $N\times N$ self-attention matrix is not strictly necessary.
Based on the aforementioned considerations, we have developed a novel, low-complexity Unified Attention Module (UAM) to tackle channel decoding tasks involving multiple code types, lengths, and rates. We introduced learnable memory components, $\textbf{{A}}_{l}$ (attention) and $\textbf{{V}}_{l}$ (values), which play the role of learning features from the entire training dataset. The mathematical expression is given by
$UniAttn\left({{\textbf{{A}}_{l}},{\textbf{{V}}_{l}}}\right)=Softmax\left({{\textbf{{A}}_{l}}}\right)\times\left({{\textbf{{V}}_{l}^{T}}\textbf{{X}}}\right)$ (11)
where $\textbf{{X}}\in\mathbb{R}^{N\times d_{k}}$ is the input of attention layer, $\textbf{{A}}_{l},\textbf{{V}}_{l}\in\mathbb{R}^{N\times d_{l}}$ are trainable parameters, and $d_{l}$ is an adjustable hyperparameter.
In order to gain a better understanding of the proposed UAM, we rewrite equation (2) as follows
$Attention\left(\textbf{{W}}^{\textit{Q}},\textbf{{W}}^{\textit{K}},\textbf{{W}}^{\textit{V}}\right)=Softmax\left(\frac{\left(\textbf{{XW}}^{\textit{Q}}\right)\times{\left(\textbf{{XW}}^{\textit{K}}\right)}^{T}}{\sqrt{d_{k}}}\right)\times\left(\textbf{{XW}}^{\textit{V}}\right)$ (12)
where $\textbf{{W}}^{\textit{Q}},\textbf{{W}}^{\textit{K}},\textbf{{W}}^{\textit{V}}\in\mathbb{R}^{d_{k}\times d_{k}}$ are the projection matrices.
Figure 5: Architecture of the vanilla multi-head self-attention (A) and the proposed multi-head unified-attention (B).
Algorithm 1 Pseudocode for Multi-Head Unified-Attention
1 Input: X  # $shape=(B,N,d_{h})$
2 Trainable Parameters: $\textbf{{A}}_{l}$, $\textbf{{V}}_{l}$
3 Hyper Parameters: $H$, number of heads; $B$, batch size
4 Hyper Parameters: $d_{k}$, embedding size; $d_{h}=H\times d_{k}$
5 $\textbf{{X}}=\textbf{{X}}.view(B,N,H,d_{k}).transpose(1,2)$
6 $\textbf{{V}}=\textbf{{V}}_{l}^{T}\times\textbf{{X}}$  # $shape=(B,H,d_{l},d_{k})$
7 $\textbf{{A}}=Softmax(\textbf{{A}}_{l},dim=-1)$  # $shape=(B,1,N,d_{l})$
[…]
9 $\textbf{{Out}}=\textbf{{Out}}.transpose(1,2).view(B,N,d_{h})$
10 $\textbf{{Out}}=\textbf{{W}}^{O}(\textbf{{Out}})$
11 Output: Out  # $shape=(B,N,d_{h})$
Comparing equations (12) and (11), it is evident that the computational complexity of the vanilla self-attention module and the proposed UAM is $O\left(d_{k}\times N^{2}\right)$ and $O\left(d_{k}\times d_{l}\times N\right)$, respectively. Here, note that we omit the Q and K projections in the vanilla Transformer which results in further parameter savings. In fact, as observed in the subsequent experiments, a small value of $d_{l}$ is sufficient to achieve a good bit error rate. Therefore, the proposed UAM is more suitable for large-scale inputs and is more accommodating to long channel codes compared to vanilla self-attention module. The proposed multi-head unified-attention and corresponding Python-style pseudocode are illustrated in Figure 5 and Algorithm 1, respectively. As depicted in Figure 5, thanks to the sparsity of the attention scores, multiple heads are able to share the same attention $\textbf{{A}}_{l}$, resulting in a more cohesive decoding architecture and further reducing the complexity of the neural network. This approach contrasts with vanilla self-attention mechanisms, in which each head independently computes its scores using separate learned weights.
3.2 Sparse Masked Attention
To detect and correct errors, a decoder needs to examine and contrast the received codeword using parity check equations, where a non-zero syndrome signals the presence of channel errors. In the Transformer architecture, attention is sparse as each parity bit does not correlate with every information bit. To address this, we propose enhancing the attention model with insights from human intelligence, which can clarify the dependencies between information and parity bits. By incorporating these dependencies, the model can more rapidly and accurately grasp the underlying principles of channel decoding, leading to enhanced performance.
Specifically, for any linear block code defined by a parity-check matrix $\textbf{H}\in\mathbb{F}_{2}^{(n-k)\times n}$, we define a mapping function $M(\textbf{H}):\{0,1\}^{(n-k)\times n}\to\{-\infty,0\}^{N\times d_{l}}$. This function operates on H and produces a mask that is applied to the attention memory $\textbf{{A}}_{l}$, such that
$UniAttn\left(\boldsymbol{A}_{l},\boldsymbol{V}_{l}\right)=Softmax\left(\boldsymbol{A}_{l}+M(\textbf{H})\right)\times\left(\boldsymbol{V}_{l}^{T}\boldsymbol{X}\right)$ (13)
Let $\textbf{{A}}_{l}$ be a matrix in $\mathbb{R}^{N\times d_{l}}$, where $N$ is defined as $2n-k$ and $d_{l}$ as $n-k$. We naturally proceed to define the density of H as
$H_{d}=\frac{sum({\textbf{H}})}{(n-k)\times n}$ (14)
We then construct the extended parity-check matrix $\overline{\textbf{H}}$ as follows
$\overline{\textbf{H}}=\begin{bmatrix}{\textbf{H}}^{T}\\ {\textbf{{J}}}_{n-k}\end{bmatrix}$ (15)
where ${\textbf{{J}}}_{n-k}$ is a randomly generated sparse matrix of size $(n-k)\times(n-k)$ with a density of $H_{d}$. The generation of the attention mask, therefore, becomes a direct process and is formulated as $M(\overline{\textbf{H}})$.
Figure 6: We present the corresponding matrices $\textbf{H}^{T}$, $\overline{\textbf{H}}$ and the Tanner graph for the Hamming (7,4) code.
For enhanced clarity, Figure 6 offers a visual depiction of matrices $\textbf{H}^{T}$, $\overline{\textbf{H}}$ and the Tanner graph, utilizing the Hamming (7,4) code as an illustrative example. Here, matrix H is defined as
$\textbf{H}=\begin{bmatrix}1&1&1&0&1&0&0\\ 1&0&1&1&0&1&0\\ 0&1&1&1&0&0&1\end{bmatrix}$ (16)
Algorithm 2 Pseudocode for Sparse Mask Construction
1 Input: H  # $shape=(n-k,n)$
2 $H_{d}=\frac{sum(\textbf{H})}{(n-k)\times n}$
3 $\overline{\textbf{H}}=zeros(2n-k,n-k)$
4 $\overline{\textbf{H}}[0:n,0:n-k]=\textbf{H}.transpose(0,1)$
5 $\overline{\textbf{H}}[n:2n-k,0:n-k]=rand(n-k,n-k)<H_{d}$
6 Output: ()(¬$(-\infty)\cdot(\neg\mathrm{~{}\overline{\textbf{H}}})$
This illustration demonstrates that the application of the mask results in sparser attention, a phenomenon particularly pronounced in low-density parity-check codes. Most importantly, the complexity of the UAM is now reduced to the density of the code $O(H_{d}\times d_{k}\times d_{l}\times N)$. Our experimental results indicate that the use of this mask has successfully reduced the computational complexity of the attention mechanism by an average of $83\%$. Furthermore, we provide a summary of the Python-style sparse mask construction within Algorithm 2.
3.3 Architecture and Training
The architecture of the proposed unified error correction code Transformer is illustrated in Figure 7. The primary innovation in our model involves substituting the vanilla attention mechanism with a multi-head unified-attention module. This modification not only decreases computational complexity but also empowers the model to discern distinctive features among diverse codewords, thereby facilitating the establishment of a unified decoding architecture. We incorporate advanced pre-training strategies to optimize the model’s error correction performance. Furthermore, we have designed standardization units to standardize the lengths of codewords and syndromes across diverse datasets, ensuring the model’s compatibility with various code types, lengths, and rates.
Figure 7: Illustration of the proposed unified error correction code Transformer.
In Figure 7, the output y from the AWGN channel passes through the pre-processing module to yield $N=2n-k$ input elements for the neural network. The decoder is constructed by concatenating $L$ Transformer encoders, which consist of unified-attention and feed-forward layers, with normalization layers in between. The output module is characterized by two fully connected layers. The initial layer compresses the multi-head output into a one-dimensional vector of size $2n-k$, with the subsequent layer transforming it into an $n$-dimensional vector representing the softly decoded noise. Finally, the post-processing module acts on the estimated noise to derive the transmitted original codeword.
We employ binary cross-entropy as the loss function with the objective of learning to predict the multiplicative noise z. The corresponding target binary multiplicative noise is denoted as $=bin(sign(\textbf{{z}}))$ z ) ). Therefore, the loss calculation for a received individual codeword y is given by
$loss=-\sum_{i=1}^{n}{\tilde{z}}_{i}log(f_{\theta}(y))+(1-{\tilde{z}}_{i})log(1-f_{\theta}(y))$ (17)
[…]
$\widetilde{\textbf{{y}}}=\left[~{}\left[\lvert\textbf{{y}}\rvert,~{}\textbf{0}_{c}\right],~{}\left[s(\textbf{{y}}),~{}\textbf{0}_{s}\right]~{}\right]$ (18)
where $\textbf{0}_{c}$ and $\textbf{0}_{s}$ represent zero vectors with sizes $N_{max}-size(\textbf{{y}})$ and $S_{max}-size(s(\textbf{{y}}))$, respectively. Through this unit, the Transformer neural decoder can accommodate any codeword within the maximum length, allowing for mixed training in a unified manner.
In the training phase, we conducted 1000 epochs, each consisting of 1024 minibatches, which in turn contained 512 samples each. Samples for each minibatch were randomly selected from a pool of systematic Polar[28], LDPC, and BCH codewords. The learning rate was initialized at $10^{-3}$, and a cosine decay scheduler was applied without warmup [29], gradually reducing it to $10^{-6}$ by the end of the training process. We employ the Adam optimizer to dynamically adjust the learning rate, leveraging its efficacy in managing diverse data types and model architectures [30]. The AWGN noise introduced to the links had its signal-to-noise ratio (SNR) randomly chosen between 4 and 6 for each batch. In the pre-training phase, we conducted 1000 epochs while maintaining consistency with all other parameters from the training phase. The word embeddings acquired during the pre-training stage served as the initial values for the subsequent training phase. The training and experiments were performed on NVIDIA GeForce RTX 4090 24GB GPUs.
4 Experiments
We conducted experiments on linear block codes, specifically Polar, LDPC, and BCH, to assess the effectiveness of the proposed architecture. The parity-check matrices utilized in the paper were sourced from [31]. We benchmark our approach against the most recent state-of-the-art performance reported in [22], where each codeword parameter was trained individually and also contrast it with the legacy successive cancellation list decoding algorithm for Polar codes [32], as well as the normalized min-sum decoding algorithm for LDPC codes [10] [33]. The results are presented in the form of bit error rate (BER) and block error rate (BLER) at various normalized SNR (i.e. $E_{b}/N_{0}$) values, with at least $10^{6}$ random codewords tested for each. The terms CBER and CBLER are hereby defined to denote the comprehensive bit error rate and block error rate, respectively, for the entirety of codewords utilized, spanning Polar, LDPC, and BCH codes.
The proposed unified error correction code Transformer is defined by the following hyperparameters: the number of Transformer encoder layers $L$, the dimensions of word embeddings $d_{k}$, and the dimensions $d_{l}$ of the trainable memory $\textbf{{A}}_{l}$ and $\textbf{{V}}_{l}$ along with the number of unified attention heads $H$. We present the performance of our framework for four different architectures with $L\in\{3,6\},d_{k}\in\{16,32\}$, while all other hyperparameters are set to their default values: $d_{l}=64,H=8$. The hyperparameters in [22] are set as follows: $L=6,d_{k}=32,H=8$. Additionally, for Polar SCL decoding, the list size is set to 8, and for LDPC NMS decoding, the maximum number of iterations is set to 15 with a normalized factor of 0.875.
Table 1 presents the simulation results, specifically showing the negative logarithm base 10 of the BER at $E_{b}/N_{0}$ values of 4, 5, and 6 dB. It is noted that higher values in this context are indicative of better performance. The best performance for each SNR is highlighted in bold and italics. We provide additional BER plots for Polar and LDPC codes in Figure 8 and Figure 9, respectively. Experimental results indicate that our approach achieves comparable performance to models trained individually as reported in [22], and outperforms both the SCL algorithm for Polar codes and the NMS algorithm for LDPC codes. Our proposed unified code-agnostic Transformer effectively decodes a wide range of code types, lengths, and rates. Especially for short LDPC codes, traditional NMS algorithms are constrained by a smaller girth, leading to poor performance because LLR information cannot propagate effectively during the iteration process. In contrast, our approach sidesteps this issue by eliminating iteration, thereby offering a practical solution for short LDPC code applications, such as those for future 6G mobile communication control channel coding schemes. Furthermore, the performance gains observed with smaller dimensions $d_{l}=64\ll N=192$ confirm that the proposed decoding architecture substantially reduces computational complexity.
Table 1: A comparison of the negative logarithm base 10 of the BER
Figure 8: BER for Polar codes under $L=6,d_{k}=32$.
Figure 9: BER for LDPC codes under $L=6,d_{k}=32$.
5 Analysis
We delve into the influence of Transformer encoder layers on overall performance, examining both the role of the dimension parameter $d_{k}$ and the efficacy of sparse masked attention mechanisms. Furthermore, we conduct a comparative analysis of the computational complexity inherent in our methodology versus that of alternative approaches.
5.1 Impact of the Transformer Encoder Layers and the Dimensions $d_{k}$
Figure 10 illustrates the CBER performance across various Transformer encoder layers and the dimensions $d_{k}$. It can be observed that the number of Transformer encoder layers exerts a greater impact on performance than the dimension $d_{k}$. This increase in depth effectively enlarges the parameter search space, potentially enhancing the model’s nonlinear fitting capability and enabling more complex expression.
The results suggest that a deeper Transformer architecture is more capable of capturing intricate patterns and dependencies within the data, which is crucial for improving error correction performance. By adding more layers, the model can perform multiple levels of abstraction and refinement, allowing it to better understand and correct errors. This is especially crucial in the context of wireless communication systems, which are characterized by complexity and variability that simple models might not adequately capture.
In essence, the findings emphasize the importance of model depth in Transformer architectures for error correction tasks. This insight can guide the design of more effective models by prioritizing the depth of the encoder layers over the dimension of $d_{k}$, leading to better performance in practical applications.
Figure 10: CBER performance across various Transformer encoder layers and the dimensions $d_{k}$.
5.2 Impact of the Sparse Masked Attention
We illustrate the impact of sparse masked attention on loss, CBER, and CBLER in Fig. 11. Specifically, sparse masked attention reduces the loss by $45\%$ and leads to faster convergence during training. Furthermore, this approach significantly improves error correction performance: the CBER shows an improvement of approximately $0.4$ dB at a bit error rate of $10^{-5}$, and the CBLER improves by about $0.5$ dB at a block error rate of $10^{-4}$.
These improvements underscore the efficacy of integrating domain-specific knowledge into the attention mechanism, demonstrating substantial gains in both efficiency and accuracy. This approach not only reduces computational complexity but also enhances the model’s ability to decode complex signals, making it a promising strategy for future advancements in error correction techniques.
Figure 11: The overall improvement in loss, CBER and CBLER with sparse masked attention.
5.3 Computational Complexity Analysis
Before delving into the computational complexity, let’s first review the unified error correction code Transformer presented in Figure 7, which incorporates $L$ layers of Transformer encoders. The principal computational complexity within each layer is found in the multi-head self-attention module, which is predominantly responsible for matrix multiplication and is defined by the hyperparameters $H,d_{k}$ and $d_{l}$. Consequently, the computational complexity of the proposed architecture is given by $O\left({L\times H\times\left({{H_{d}}\times{d_{k}}\times{d_{l}}\times N+d_{k}^{2}}\right)}\right)$, where $N=2n-k$, $d_{l}\ll N$ and $H\ll N$. Thanks to the sparsity of the parity-check matrix, $H_{d}$ is usually quite small. In our experiment, $H_{d}$ is approximately 0.17, which significantly reduces the computational complexity of the neural network.
In comparison, the approach in [22], which employs the vanilla Transformer architecture, has a complexity of $O\left({L\times H\times\left({{N^{2}}\times{d_{k}}+d_{k}^{2}}\right)}\right)$. When it comes to non-learning solutions, the computational complexities of the SCL and NMS decoding algorithms are $O\left(L_{s}\times n\times log_{2}\left(n\right)\right)+O\left(L_{s}\times\left(n-1\right)\right)+k\times O\left(2\times L_{s}\times log_{2}\left(2\times L_{s}\right)\right)$ and $I_{max}\times O\left(2\times n\times d_{\nu}+M\times\left(2\times d_{c}-1\right)\right)+I_{max}\times O\left(M\times d_{c}\right)$, respectively [34]. Here, $L_{s}$ is the list size, $I_{max}$ is the maximum number of iterations, $M$ is the size of the check matrix, $d_{v}$ is the number of variable node, and $d_{c}$ is the number of parity check node.
Compared to the approach in [22], while the proposed method excels in decoding versatility and computational complexity, it may fall short in terms of memory efficiency, power consumption, and computational resource utilization when compared to non-learning solutions. This inefficiency could limit its potential for deployment. To address these issues, future implementations could leverage techniques such as parameter sharing, low-rank factorization, and quantization, aiming to reduce complexity and enhance efficiency [35, 36].
6 Conclusions
We introduced a novel, unified code-agnostic Transformer decoding architecture, designed to integrate various linear block codes seamlessly and offer superior decoding performance. Our approach leverages the self-attention mechanism inherent in Transformer models, enabling effective handling of sequential data and capturing long-range dependencies critical for decoding tasks. Through rigorous analysis and extensive experimental validation, our results indicate that the unified Transformer decoding architecture can significantly advance the state-of-the-art in decoding performance, potentially setting new benchmarks for future research and practical applications.
In conclusion, the unified Transformer decoding architecture presented in this paper signifies a significant step forward in the field of error correction coding. It paves the way for the development of more adaptable, efficient, and robust decoding solutions, poised to meet the stringent requirements of 6G and beyond. As the field progresses, we anticipate that the continued integration of AI technologies will unlock new possibilities, driving innovation in wireless communication systems. Future work can aim to further reduce hardware implementation complexity and enhance computational efficiency using techniques such as quantization and compression of neural network models.

### v2

@@
 For instance, the Successive Cancellation (SC) decoding algorithm [4] for Polar codes uses a sequential method that traverses a decision tree to obtain the decoding result.
-This approach simplifies the hardware implementation of Polar codes by pruning binary trees [9].
+This approach simplifies the hardware implementation of Polar codes by pruning binary trees [9, 10, 11].
 In contrast, the Sum-Product (SP) decoding algorithm [12] for LDPC codes uses iterative message passing, allowing for parallel data processing.
@@
 This may lead to significant engineering effort and inefficient use of hardware resources.
-In an effort to reduce hardware resource consumption and establish a quasi-unified decoding framework, three methods have been explored: resource sharing, noise estimation, and Artificial Intelligence (AI)-driven approaches.
+In an effort to reduce hardware resource consumption and establish a quasi-unified decoding framework, three methods have been explored: resource sharing, Ordered-Statistics Decoding (OSD), and Artificial Intelligence (AI)-driven approaches.
 Resource sharing involves time-sharing storage and computing units.
-Noise estimation methods decode various linear block codes by estimating the noise of received signals.
+Ordered-Statistics Decoding leverages the reliability ordering of received symbols to efficiently decode various linear block codes by systematically evaluating the most likely codewords.
 Meanwhile, AI-driven methods offer a convenient way to unfold traditional algorithms like Belief Propagation (BP).
@@
 However, due to the distinct decoding algorithms for each code type, achieving a truly unified architecture with such decoders is generally challenging.
-Second, noise estimation methods were first proposed in [15] as the Guessing Random Additive Noise Decoding (GRAND) algorithm, used for Maximum Likelihood (ML) decoding of arbitrary codebooks in discrete channels, with or without memory [16, 17].
-The algorithm operates by rank-ordering noise sequences based on their likelihood and subtracting them from the received signal to identify codebook members [16].
-GRAND is shown to be capacity-achieving with random codebooks.
-However, this work faces the challenge of excessive guessing required by the GRAND algorithm, particularly for high logistic weights, which can lead to computational inefficiency.
-For instance, when the logistic weight equals $128$, up to $5.33\times 10^{7}$ queries may be necessary to complete a single decoding process [17].
+Second, ordered-statistics decoding methods, initially explored in [17, 18, 19], provide a near-Maximum Likelihood (ML) decoding approach for linear block codes by exploiting the reliability ordering of received symbols.
+The algorithm operates by sorting the received signal positions based on their reliability, permuting the code’s generator matrix accordingly, and systematically evaluating the most likely codewords through a limited search.
+OSD delivers near-ML performance with significantly less complexity than exhaustive ML decoding, making it well-suited for short to moderate-length codes.
+However, its drawback lies in the computational complexity, which grows exponentially with the search order (e.g., order-$l$).
+For instance, an order-4 OSD may require evaluating thousands of test codewords, potentially leading to increased latency and resource demands.
 Third, in recent years, AI-driven decoders have gained attention as an alternative.
@@
 In contrast, model-free approaches offer a more flexible and scalable solution.
-Unlike model-based methods, which depend on predefined mathematical models, model-free approaches do not assume prior knowledge of the encoder or decoder.
+Unlike model-based methods, which depend on predefined mathematical models, model-free approaches assume no prior knowledge of the encoder or decoder.
 Instead, they rely on large amounts of labeled data to learn the decoding process directly, bypassing the limitations of conventional algorithms [22].
-For example, inspired by the success of Transformers in Natural Language Processing (NLP) [21], a Transformer-based, model-free neural decoder called the Error Correction Code Transformer (ECCT) is proposed [22].
-ECCT leverages the Transformer’s self-attention mechanism, where multiple attention heads work in parallel to effectively manage sequential data and capture long-range dependencies, to establish relationships between the input Log-Likelihood Ratios (LLRs) in the decoder.
+For instance, inspired by the success of Transformers in Natural Language Processing (NLP) [23], a model-free neural decoder called the Error Correction Code Transformer (ECCT) was proposed [24].
+ECCT employs the Transformer’s self-attention mechanism, where multiple attention heads operate in parallel to process sequential data and capture long-range dependencies, effectively linking input Log-Likelihood Ratios (LLRs) to the decoder’s output.
 The inherent connection between Transformers and decoders lies in their shared ability to translate sequences—whether from one language to another or from LLRs back to the original information bits.
-However, despite offering greater flexibility, model-free approaches like ECCT still require frequent tuning of model parameters for each code type, length, and rate.
-This increases training costs and hardware overhead, making the decoding architecture non-unified.
+However, ECCT’s flexibility comes at a cost, requiring frequent parameter adjustments for different code types, lengths, and rates, which increases training complexity and hardware overhead, hindering a unified decoding architecture.
+Building on this, [25] introduced a Foundation Model for Error Correction Codes (FECCT) to improve ECCT’s generalization to unseen codewords.
+Yet, FECCT currently faces the challenge of high computational complexity in its self-attention mechanism, further complicating its practical implementation.
 1.2 Contributions
@@
 The unified attention compresses the structural information of different codewords through learning and shares parameters across various heads in the Transformer, resulting in a more cohesive decoding architecture.
-Additionally, we eliminated the need for iterations through the unified attention mechanism, thereby avoiding the detrimental effects of short cycles in Tanner graphs on LDPC belief propagation decoders.
-As a result, we have unlocked the full potential of LDPC codes, extending their applicability to short code scenarios.
   4. 4.
@@
 This results in a significant reduction in complexity from the traditional $O(N^{2})$ to a more efficient $O(N)$, where $N$ represents the sequence length.
-Additionally, by employing our sparse mask, we further achieve an $83\%$ reduction in computational complexity.
+Additionally, by employing our sparse mask, we further achieve an $86\%$ reduction in computational complexity.
 This efficiency gain enables faster processing and lower power consumption, making the architecture suitable for real-time applications in next-generation wireless communication systems.
@@
 Finally, Section VI concludes.
-Notations: Lowercase and uppercase bold letters represent vectors and matrices, respectively.
-GF(2) denotes the binary Galois field.
-$\mathbb{R}$ represents real numbers.
-$[\cdot]^{T}$ stands for transpose, while $\mathcal{N}(\mu,{\sigma}^{2})$ signifies a Gaussian distribution with mean $\mu$ and variance ${\sigma}^{2}$.
-We define $bipolar(\cdot)$ to map binary to bipolar signals, and $bin(\cdot)$ represents its inverse mapping, where $bipolar(x)$ is defined as $1-2x$, and $x$ takes values in the set $\{0,1\}$.
-$size(\cdot)$ represents the number of elements in a matrix or vector.
-The sign function is denoted as $sign(\cdot)$.
+Notations: Bold lowercase and uppercase letters denote vectors (e.g., $\mathbf{v}$) and matrices (e.g., $\mathbf{M}$), respectively, while scalars use regular letters (e.g., $x$).
+$\mathbb{F}_{2}$ represents the Galois field of order 2, and $\mathbb{R}$, $\mathbb{N}$ denote the real and natural numbers.
+The transpose is indicated by $[\cdot]^{T}$, and $\mathcal{N}(\mu,\sigma^{2})$ denotes a Gaussian distribution with mean $\mu$ and variance $\sigma^{2}$.
+The function $\text{bipolar}(x)=1-2x$ maps $x\in{0,1}$ to bipolar values ${1,-1}$, with its inverse $\text{bin}(x)=\frac{1-x}{2}$ mapping ${1,-1}$ back to ${0,1}$.
+The operator $\text{size}(\cdot)$ denotes the number of elements in a vector or matrix, and $\text{sign}(\cdot)$ represents the sign function, defined as $\text{sign}(x)=1$ if $x>0$, 00 if $x=0$, and $-1$ if $x<0$.
+For a linear block code with code length $n$ and information bit length $k$, we denote it as $(n,k)$.
 2 Background
@@
 2.1 Vanilla Transformers
-The vanilla Transformer functions as a sequence-to-sequence model comprising an encoder and a decoder, both built from a series of $L$ identical blocks.
-Each encoder block primarily integrates a multi-head self-attention module and a position-wise feed-forward network (FFN).
-To construct a deeper model and prevent gradient degradation, residual connections[23] are employed between the self-attention module and the FFN.
-Layer normalization[24] is applied to the inputs of each module, enhancing network stability and convergence speed.
-The overall architecture of the single-layer vanilla Transformer encoder is illustrated in Figure 1.
-Figure 1: Overall architecture of the single-layer vanilla Transformer.
-The following subsections will introduce the key modules within the vanilla Transformer.
-2.1.1 Word Embedding
-Word embedding is an important term in natural language processing, used to represent words for text analysis in the form of real-valued vectors.
-It enhances the computer’s ability to better understand text-based content.
-In this paper, word embeddings serve as reliability encoders that map the received one-dimensional signal to higher dimensions.
-It can be expressed as
-$Embedding\left(\textbf{{X}}\right)=\textbf{{X}}\odot\textbf{{E}}$ (1)
-where $\textbf{{X}}\in\mathbb{R}^{N\times 1}$ is the received signal with length $N$ and dimension one, $\textbf{{E}}\in\mathbb{R}^{N\times d_{k}}$ represents the trainable embedding matrix with dimensions of $d_{k}$, and $\odot$ denotes the Hadamard product.
-2.1.2 Multi-Head Self-Attention
-The vanilla Transformer establishes an attention mechanism based on the three Query-Key-Value (Q-K-V) matrices.
-It utilizes the Q and K matrices to establish correlations between different elements within the sequence and weights the V matrix accordingly.
-The scaled dot-product form of attention is given by
-$Attention\left(\textbf{{Q}},\textbf{{K}},\textbf{{V}}\right)=Softmax\left(\frac{\textbf{{QK}}^{T}}{\sqrt{d_{k}}}\right)\textbf{{V}}$ (2)
-$Softmax(\textbf{{z}})_{i}=\frac{e^{z_{i}}}{\sum_{j=1}^{n}{e^{z_{j}}}}~{}for~{}i=1,2,\cdots,n$ (3)
-where $\textbf{{Q}}\in\mathbb{R}^{N\times d_{k}}$, $\textbf{{K}}\in\mathbb{R}^{M\times d_{k}}$, $\textbf{{V}}\in\mathbb{R}^{M\times d_{v}}$ represent tensors with lengths of $N$ (for queries and keys) and $M$ (for values), and dimensions of $d_{k}$ (for queries and keys) and $d_{v}$ (for values), respectively.
-The $Softmax\left(\cdot\right)$ function normalizes inputs into the range of 0 to 1, and to alleviate its gradient degradation issue, the dot product of queries and keys is divided by $\sqrt{d_{k}}$.
-The study in[21] enhanced the attention layer by introducing a technique known as multi-head attention, which involves $H$ different learned projection sets.
-This enhancement benefits the attention layer’s performance in two significant ways: it widens the model’s capacity to concentrate on various positions and grants the attention layer multiple representation subspaces.
-The multi-head attention can be written as
-$MultiHeadAttn\left(\textbf{{Q}},\textbf{{K}},\textbf{{V}}\right)=Concat\left(head_{1},head_{2},\cdots,head_{H}\right)\textbf{{W}}^{\textit{O}}$ (4)
-where $head_{i}=Attention(\textbf{{QW}}_{i}^{\textit{Q}},\textbf{{KW}}_{i}^{\textit{K}},\textbf{{VW}}_{i}^{\textit{V}})$, and $\textbf{{W}}_{i}^{\textit{Q}}\in\mathbb{R}^{d_{k}\times d_{k}}$, $\textbf{{W}}_{i}^{\textit{K}}\in\mathbb{R}^{d_{k}\times d_{k}}$, $\textbf{{W}}_{i}^{\textit{V}}\in\mathbb{R}^{d_{v}\times d_{v}}$ are trainable parameters, $\textbf{{W}}^{\textit{O}}\in\mathbb{R}^{Hd_{v}\times Hd_{v}}$ is a weight matrix that was trained jointly with the model, serving to condense the outputs of different heads down into a single matrix.
+The vanilla Transformer, a sequence-to-sequence model, comprises an encoder and decoder, each with $L$ identical blocks.
+Each block integrates multi-head self-attention (MHA), a position-wise feed-forward network (FFN), residual connections [26] to mitigate gradient issues, and layer normalization [27] for stability and faster convergence, as shown in Figure 1.
+Figure 1: Single-layer vanilla Transformer encoder architecture.
+The process begins with embedding a one-dimensional input signal $\textbf{{x}}\in\mathbb{R}^{N}$ into a higher-dimensional space:
+$\textbf{{X}}_{e}=\text{Embedding}(\textbf{{x}})=\text{diag}(\textbf{{x}})\cdot\textbf{{W}}$ (1)
+where $\text{diag}(\textbf{{x}})\in\mathbb{R}^{N\times N}$ is a diagonal matrix with x on the diagonal, and $\textbf{{W}}\in\mathbb{R}^{N\times d_{k}}$ is a trainable weight matrix, producing $\textbf{{X}}_{e}\in\mathbb{R}^{N\times d_{k}}$ with $d_{k}$-dimensional embeddings for each position.
+This embedding feeds into the MHA mechanism, which correlates sequence elements through:
+$\text{Attn}\left(\textbf{{Q}},\textbf{{K}},\textbf{{V}}\right)=\text{Softmax}\left(\frac{\textbf{{Q}}\cdot\textbf{{K}}^{T}}{\sqrt{d_{k}}}\right)\cdot\textbf{{V}}$ (2)
+extended to multiple heads as:
+$\text{MHA}\left(\textbf{{Q}},\textbf{{K}},\textbf{{V}}\right)=\text{Concat}\left(\text{head}_{1},\cdots,\text{head}_{H}\right)\cdot\textbf{{W}}^{\textit{O}}$ (3)
+where $\text{head}_{i}=\text{Attn}(\textbf{{QW}}_{i}^{Q},\textbf{{KW}}_{i}^{K},\textbf{{VW}}_{i}^{V})$, with $\textbf{{Q}},\textbf{{K}},\textbf{{V}}\in\mathbb{R}^{N\times d_{k}}$ set to $\textbf{{X}}_{e}$ for self-attention, using trainable weights $\textbf{{W}}_{i}^{Q},\textbf{{W}}_{i}^{K},\textbf{{W}}_{i}^{V}\in\mathbb{R}^{d_{k}\times d_{k}}$ and $\textbf{{W}}^{O}\in\mathbb{R}^{Hd_{k}\times Hd_{k}}$, as depicted in Figure 2.
 Figure 2: Architecture of multi-head self-attention.
-[…]
-2.1.3 Position-Wise Feed-Forward Network
-The outputs of the self-attention layer are fed into the position-wise FFN layer.
-The position-wise FFN layer is essentially a fully connected network that operates uniformly on each position, and the non-linear activation function within it enables the Transformer to learn more complex feature representations.
-The mathematical expression for the FFN is as follows
-$FFN\left(\textbf{{X}}\right)=\textbf{{W}}_{2}\times ReLU\left(\textbf{{W}}_{1}\times\textbf{{X}}+\textbf{{b}}_{1}\right)+\textbf{{b}}_{2}$ (5)
-$ReLU\left(x\right)=max\left(0,x\right)$ (6)
-where X is the output of previous self-attention layer, and $\textbf{{W}}_{1}\in\mathbb{R}^{d_{f}\times d_{h}}$, $\textbf{{W}}_{2}\in\mathbb{R}^{d_{h}\times d_{f}}$, $\textbf{{b}}_{1}\in\mathbb{R}^{d_{f}}$, $\textbf{{b}}_{2}\in\mathbb{R}^{d_{h}}$ are trainable parameters, $d_{h}=H\times d_{k}$ is the dimensions of queries, keys and values, $ReLU\left(\cdot\right)$ represents the activation function.
-In general, to achieve better performance, $d_{f}$ is set to be greater than $d_{h}$.
-2.1.4 Residual Connection and Normalization
-The multi-head self-attention layer and the position-wise FFN layer are connected through residual network[23] and layer normalization[24].
-The overall architecture of each vanilla Transformer encoder can be written as
-$\textbf{{X}}_{1}$ $=LayerNorm\left(MultiHeadAttn(\textbf{{X}})+\textbf{{X}}\right)$ (7)
-$\textbf{{X}}_{2}$ $=LayerNorm\left(FFN(\textbf{{X}}_{1})+\textbf{{X}}_{1}\right)$
-where $LayerNorm\left(\cdot\right)$ denotes the layer normalization module.
+Next, the MHA output is processed by the FFN, applied uniformly across positions:
+$\text{FFN}\left(\textbf{{X}}\right)=\textbf{{W}}_{2}\cdot\text{ReLU}\left(\textbf{{W}}_{1}\cdot\textbf{{X}}+\textbf{{b}}_{1}\right)+\textbf{{b}}_{2}$ (4)
+where $\textbf{{W}}_{1}\in\mathbb{R}^{d_{f}\times d_{h}}$, $\textbf{{W}}_{2}\in\mathbb{R}^{d_{h}\times d_{f}}$, $\textbf{{b}}_{1}$, $\textbf{{b}}_{2}$ are trainable parameters, $d_{h}=H\cdot d_{k}$ and $\text{ReLU}(x)=\max(0,x)$ enables complex feature learning, with $d_{f}>d_{h}$ for enhanced performance.
+Finally, residual connections and layer normalization link the MHA and FFN layers:
+$\textbf{{X}}_{1}$ $=\text{MHA}\left(\text{LayerNorm}\left(\textbf{{X}}_{e}\right)\right)+\textbf{{X}}_{e}$ (5)
+$\textbf{{X}}_{2}$ $=\text{FFN}\left(\text{LayerNorm}\left(\textbf{{X}}_{1}\right)\right)+\textbf{{X}}_{1}$
 2.2 Forward Error Correction Codes
 In information theory, FEC is a technique designed to reduce the bit error rate (BER) when transmitting data over a noisy channel.
-We establish a standard linear block code utilizing a generator matrix G with dimensions $k\times n$ and a parity-check matrix H with dimensions $(n-k)\times n$.
-These matrices are specifically designed to satisfy $\textbf{G}\textbf{H}^{T}=\textbf{0}$ over the binary Galois field (GF).
-We will now briefly discuss the encoder.
-The message vector $\textbf{{m}}\in\{0,1\}^{k}$ multiplied with the generator matrix G over the GF(2) results in a codeword vector $\textbf{{x}}\in\{0,1\}^{n}$ satisfying $\textbf{H}\textbf{{x}}=\textbf{0}$.
-The encoded codeword x is modulated by Binary Phase Shift Keying (BPSK) to $\textbf{{x}}_{s}$, where $BPSK(x)=1-2x$.
-$\textbf{{x}}_{s}$ is transmitted through binary-input discrete memoryless channels (B-DMCs, e.g.
-an Additive White Gaussian Noise channel), resulting in a channel output $\textbf{{y}}=\textbf{{x}}_{s}+\textbf{{n}}$, where n is Gaussian noise with zero mean and variance $\sigma^{2}$.
+We adopt a linear block code defined by a generator matrix $\textbf{G}\in\mathbb{R}^{k\times n}$ and a parity-check matrix $\textbf{H}\in\mathbb{R}^{(n-k)\times n}$, satisfying $\textbf{G}\cdot\textbf{H}^{T}=\textbf{0}$ over $\mathbb{F}_{2}$.
+[…]
 Figure 3: Architecture of the communication system.
-The goal of the decoder is to recover the transmitted message m from the received signal y.
-The overall architecture of the communication system including both encoding and decoding is illustrated in Figure 3.
-The objective of our work is to design and training of the parameterized decoding function $\textit{f}_{\theta}$.
 Figure 4: Decoder framework with pre- and post-processing.
-To build a model-free Transformer decoding model, we draw inspiration from the pre- and post-processing methods proposed in [25], which have been demonstrated not to result in performance loss concerning bit error rate or mean squared error (MSE) metrics, as depicted in Figure 4.
-By training with an all-zero codeword to learn the noise, it is possible to avoid the overfitting problem that arises from the exponential growth of the codeword space with code length.
+The decoder aims to recover m from y, as depicted in Figure 3.
+Our work focuses on designing and training a parameterized decoding function $f_{\theta}$.
+To develop a model-free Transformer-based decoder, we leverage pre- and post-processing techniques from [28], preserving BER and Mean Squared Error (MSE) performance (see Figure 4).
+Training with an all-zero codeword mitigates overfitting due to exponential codeword space growth.
 And the B-DMCs is modeled by (6) which is an equivalent statistical mode that differs from the true physical one [29].
@@
 where z is a random multiplicative noise independent of ${\textbf{{x}}_{s}}$.
-In the pre-processing step, the absolute values of received signal and the syndromes are concatenated as the input to the neural network, which can be written as
-$=\left[\lvert\textbf{{y}}\rvert,s\left(\textbf{{y}}\right)\right]$ y ) ] (9)
+In the pre-processing step, the absolute values of received signal and the syndromes are concatenated as the input to the neural network, which can be written as:
+y~=[|y|,s($\textbf{{y}}\right)\right]$ = [ | y | , s ( y ) ] (7)
 […]
-In the post-processing step, the predicted multiplicative noise $\hat{\textbf{{z}}}$ is multiplied by the received signal y to recover x. That is, the predicted $\hat{\textbf{{x}}}$ takes the form
-$\hat{\textbf{{x}}}=bin\left(sign\left({\textbf{{y}}}\cdot{\hat{\textbf{{z}}}}\right)\right)$ (10)
+In the post-processing step, the predicted noise $\hat{\textbf{{z}}}$ is multiplied by the received signal y to recover x. That is, the predicted $\hat{\textbf{{x}}}$ takes the form:
+$\hat{\textbf{{x}}}=\text{bin}\left(\text{sign}\left({\textbf{{y}}}\cdot{\hat{\textbf{{z}}}}\right)\right)$ (8)
+2.3 Mask Generation in ECCT and FECCT
+Masks in Transformer-based error correction enhance decoding by integrating code structure into attention mechanisms.
+In ECCT [24], the mask $g(\textbf{H}):\{0,1\}^{(n-k)\times n}\to\{-\infty,0\}^{(2n-k)\times(2n-k)}$ embeds code structure using the parity-check matrix H.
+Initialized as an identity matrix, it unmasks positions of paired ones in H rows, extending the Tanner graph to two-ring connectivity.
+It is applied additively in attention: $\text{Attn}(\textbf{{Q}},\textbf{{K}},\textbf{{V}})=\text{Softmax}\left(\frac{\textbf{{Q}}\cdot\textbf{{K}}^{T}+g(\textbf{H})}{\sqrt{d_{k}}}\right)\cdot\textbf{{V}}$.
+In FECCT [25], a learned $\psi:\mathbb{N}\to\mathbb{R}$ modulates attention via the distance matrix $\mathcal{G}(\textbf{H})\in\mathbb{N}^{(2n-k)\times(2n-k)}$, applied with a Hadamard product: $\text{Attn}(\textbf{{Q}},\textbf{{K}},\textbf{{V}})=\left(\text{Softmax}\left(\frac{\textbf{{Q}}\cdot\textbf{{K}}^{T}}{\sqrt{d_{k}}}\right)\odot\psi(\mathcal{G}(\textbf{H}))\right)\cdot\textbf{{V}}$.
 3 Proposed Unified Error Correction Code Transformer
 In this section, we present the complete architecture and training process of the proposed Unified Error Correction Code Transformer (UECCT).
-3.1 Unified Attention
+3.1 Shared Unified Attention
 The vanilla self-attention module can be viewed as establishing correlations among different input features within a data sample while neglecting the intrinsic relationships between different samples.
 This limitation hinders its ability to learn the distinctiveness of features between various code types in channel decoding tasks.
-Furthermore, research in [27] suggests that the output of the softmax function is probabilistically sparse, indicating that only a few weights dominate.
+Furthermore, research in [30] suggests that the input of the softmax function is probabilistically sparse, indicating that only a few weights dominate.
 This implies that an $N\times N$ self-attention matrix is not strictly necessary.
+This sparsity suggests that a low-rank approximation could effectively capture the dominant correlations without sacrificing performance.
+Figure 5: Rank distribution of the $\textbf{{Q}}\cdot\textbf{{K}}^{T}$ matrix in the ECCT model across 6 encoder layers for various code types.
+To validate this, we trained an ECCT model (with parameters $L=6$, $H=8$, $d_{k}=32$, $d_{f}=1024$) on a variety of linear block codes, including Polar(32, 16), Polar(64, 32), Polar(128, 86), LDPC(49, 24), LDPC(121, 60), LDPC(121, 80), BCH(31, 16), BCH(63, 36), and BCH(127, 120).
+We analyzed the rank distribution of the $\textbf{{Q}}\cdot\textbf{{K}}^{T}$ matrix across 1000 batches for the 6-layer Transformer encoder.
+As shown in Figure 5, the rank ranges from 11 to 32, consistently indicating a low-rank structure across all layers and code types.
 Based on the aforementioned considerations, we have developed a novel, low-complexity Unified Attention Module (UAM) to tackle channel decoding tasks involving multiple code types, lengths, and rates.
 We introduced learnable memory components, $\textbf{{A}}_{l}$ (attention) and $\textbf{{V}}_{l}$ (values), which play the role of learning features from the entire training dataset.
-The mathematical expression is given by
-$UniAttn\left({{\textbf{{A}}_{l}},{\textbf{{V}}_{l}}}\right)=Softmax\left({{\textbf{{A}}_{l}}}\right)\times\left({{\textbf{{V}}_{l}^{T}}\textbf{{X}}}\right)$ (11)
-where $\textbf{{X}}\in\mathbb{R}^{N\times d_{k}}$ is the input of attention layer, $\textbf{{A}}_{l},\textbf{{V}}_{l}\in\mathbb{R}^{N\times d_{l}}$ are trainable parameters, and $d_{l}$ is an adjustable hyperparameter.
-In order to gain a better understanding of the proposed UAM, we rewrite equation (2) as follows
-$Attention\left(\textbf{{W}}^{\textit{Q}},\textbf{{W}}^{\textit{K}},\textbf{{W}}^{\textit{V}}\right)=Softmax\left(\frac{\left(\textbf{{XW}}^{\textit{Q}}\right)\times{\left(\textbf{{XW}}^{\textit{K}}\right)}^{T}}{\sqrt{d_{k}}}\right)\times\left(\textbf{{XW}}^{\textit{V}}\right)$ (12)
-where $\textbf{{W}}^{\textit{Q}},\textbf{{W}}^{\textit{K}},\textbf{{W}}^{\textit{V}}\in\mathbb{R}^{d_{k}\times d_{k}}$ are the projection matrices.
+The mathematical expression is given by:
+$\text{UniAttn}\left({{\textbf{{A}}_{l}},{\textbf{{V}}_{l}}}\right)=\text{Softmax}\left({{\textbf{{A}}_{l}}}\right)\cdot\left({{\textbf{{V}}_{l}^{T}}\textbf{{X}}}\right)$ (9)
+where $\textbf{{X}}\in\mathbb{R}^{N\times d_{k}}$ is the input of attention layer, $\textbf{{A}}_{l},\textbf{{V}}_{l}\in\mathbb{R}^{N\times d_{l}}$ are trainable parameters, and $d_{l}$ is an adjustable hyperparameter controlling the low-rank approximation.
+In order to gain a better understanding of the proposed UAM, we rewrite equation (2) as follows:
+$\text{Attn}\left(\textbf{{W}}^{\textit{Q}},\textbf{{W}}^{\textit{K}},\textbf{{W}}^{\textit{V}}\right)=\text{Softmax}\left(\frac{\left(\textbf{{XW}}^{\textit{Q}}\right)\cdot{\left(\textbf{{XW}}^{\textit{K}}\right)}^{T}}{\sqrt{d_{k}}}\right)\cdot\left(\textbf{{XW}}^{\textit{V}}\right)$ (10)
 Figure 6: Architecture of the vanilla multi-head self-attention (A) and the proposed multi-head unified-attention (B).
 Algorithm 1 Pseudocode for Multi-Head Unified-Attention
-1 Input: X  # $shape=(B,N,d_{h})$
+1 Input: X  # $\text{shape}=(B,N,d_{h})$
 2 Trainable Parameters: $\textbf{{A}}_{l}$, $\textbf{{V}}_{l}$
-3 Hyper Parameters: $H$, number of heads; $B$, batch size
-4 Hyper Parameters: $d_{k}$, embedding size; $d_{h}=H\times d_{k}$
-5 $\textbf{{X}}=\textbf{{X}}.view(B,N,H,d_{k}).transpose(1,2)$
-6 $\textbf{{V}}=\textbf{{V}}_{l}^{T}\times\textbf{{X}}$  # $shape=(B,H,d_{l},d_{k})$
-7 $\textbf{{A}}=Softmax(\textbf{{A}}_{l},dim=-1)$  # $shape=(B,1,N,d_{l})$
-[…]
-9 $\textbf{{Out}}=\textbf{{Out}}.transpose(1,2).view(B,N,d_{h})$
-10 $\textbf{{Out}}=\textbf{{W}}^{O}(\textbf{{Out}})$
-11 Output: Out  # $shape=(B,N,d_{h})$
-Comparing equations (12) and (11), it is evident that the computational complexity of the vanilla self-attention module and the proposed UAM is $O\left(d_{k}\times N^{2}\right)$ and $O\left(d_{k}\times d_{l}\times N\right)$, respectively.
-Here, note that we omit the Q and K projections in the vanilla Transformer which results in further parameter savings.
-In fact, as observed in the subsequent experiments, a small value of $d_{l}$ is sufficient to achieve a good bit error rate.
-Therefore, the proposed UAM is more suitable for large-scale inputs and is more accommodating to long channel codes compared to vanilla self-attention module.
+3 Hyper Parameters: number of heads $H$; batch size $B$
+4 Hyper Parameters: embedding size $d_{k}$; $d_{h}=H\times d_{k}$
+5 $\textbf{{X}}=\textbf{{X}}.\text{view}(B,N,H,d_{k}).\text{transpose}(1,2)$
+6 $\textbf{{A}}=\text{Softmax}(\textbf{{A}}_{l},\text{dim}=-1)$  # $\text{shape}=(B,1,N,d_{l})$
+7 $_{o}=\textbf{{A}}\cdot\left(\textbf{{V}}_{l}^{T}\cdot\textbf{{X}}\right)$ X )  # $\text{shape}=(B,H,N,d_{k})$
+8 $\textbf{{X}}_{o}=\textbf{{X}}_{o}.\text{transpose}(1,2).\text{view}(B,N,d_{h})$
+9 $\textbf{{X}}_{o}=\textbf{{W}}^{O}(\textbf{{X}}_{o})$
+10 Output: $\textbf{{X}}_{o}$  # $\text{shape}=(B,N,d_{h})$
+By comparing equations (9) and (10), the computational complexity of the vanilla self-attention module is $O(N^{2}\times d_{k}+N\times d_{k}^{2})$, whereas the proposed UAM reduces it to $O(N\times d_{l}\times d_{k})$.
+Importantly, the omission of the Q and K projections in the vanilla Transformer leads to significant parameter savings.
+As demonstrated in subsequent experiments, a $d_{l}$ value comparable to $n-k$ is sufficient to achieve a favorable bit error rate.
+This is supported by the fact that the maximum rank of a parity-check matrix $\textbf{H}\in\{0,1\}^{(n-k)\times n}$ is $n-k$, providing a theoretical basis for the observed performance.
+It is also important to note that both ECCT and FECCT are based on the vanilla Transformer architecture, differing only in their self-attention mask designs.
 The proposed multi-head unified-attention and corresponding Python-style pseudocode are illustrated in Figure 6 and Algorithm 1, respectively.
@@
 This approach contrasts with vanilla self-attention mechanisms, in which each head independently computes its scores using separate learned weights.
-3.2 Sparse Masked Attention
-To detect and correct errors, a decoder needs to examine and contrast the received codeword using parity check equations, where a non-zero syndrome signals the presence of channel errors.
-In the Transformer architecture, attention is sparse as each parity bit does not correlate with every information bit.
-To address this, we propose enhancing the attention model with insights from human intelligence, which can clarify the dependencies between information and parity bits.
-By incorporating these dependencies, the model can more rapidly and accurately grasp the underlying principles of channel decoding, leading to enhanced performance.
-Specifically, for any linear block code defined by a parity-check matrix $\textbf{H}\in\mathbb{F}_{2}^{(n-k)\times n}$, we define a mapping function $M(\textbf{H}):\{0,1\}^{(n-k)\times n}\to\{-\infty,0\}^{N\times d_{l}}$.
-This function operates on H and produces a mask that is applied to the attention memory $\textbf{{A}}_{l}$, such that
-$UniAttn\left(\boldsymbol{A}_{l},\boldsymbol{V}_{l}\right)=Softmax\left(\boldsymbol{A}_{l}+M(\textbf{H})\right)\times\left(\boldsymbol{V}_{l}^{T}\boldsymbol{X}\right)$ (13)
-Let $\textbf{{A}}_{l}$ be a matrix in $\mathbb{R}^{N\times d_{l}}$, where $N$ is defined as $2n-k$ and $d_{l}$ as $n-k$. We naturally proceed to define the density of H as
-$H_{d}=\frac{sum({\textbf{H}})}{(n-k)\times n}$ (14)
-We then construct the extended parity-check matrix $\overline{\textbf{H}}$ as follows
-$\overline{\textbf{H}}=\begin{bmatrix}{\textbf{H}}^{T}\\ {\textbf{{J}}}_{n-k}\end{bmatrix}$ (15)
-where ${\textbf{{J}}}_{n-k}$ is a randomly generated sparse matrix of size $(n-k)\times(n-k)$ with a density of $H_{d}$.
-The generation of the attention mask, therefore, becomes a direct process and is formulated as $M(\overline{\textbf{H}})$.
+3.2 Sparse Masked Unified Attention
+Error detection and correction in channel decoding rely on analyzing received codewords against parity-check equations, where a non-zero syndrome indicates channel errors.
+In Transformer architectures, vanilla self-attention is inherently dense, correlating all bits indiscriminately, despite parity bits not interacting with every information bit.
+To address this, we propose a sparse masked unified attention mechanism that incorporates code-specific dependencies, inspired by human intelligence in discerning bit relationships.
+This approach accelerates the model’s understanding of decoding principles, enhancing performance.
+Specifically, for any linear block code defined by a parity-check matrix $\textbf{H}\in\mathbb{F}_{2}^{(n-k)\times n}$, we define a mapping $\text{M}(\textbf{H}):\{0,1\}^{(n-k)\times n}\to\{-\infty,0\}^{N\times d_{l}}$ that generates a mask for the attention memory $\textbf{{A}}_{l}$.
+The UAM is formulated as:
+$\text{UniAttn}\left(\boldsymbol{A}_{l},\boldsymbol{V}_{l}\right)=\text{Softmax}\left(\boldsymbol{A}_{l}+\text{M}(\textbf{H})\right)\cdot\left(\boldsymbol{V}_{l}^{T}\boldsymbol{X}\right)$ (11)
+[…]
+$\left[\overline{\textbf{{y}}},\text{s}\left(\textbf{{y}}\right)\right]\cdot\overline{\textbf{H}}=\textbf{0}$ (12)
+Inspired by this, we explore a mask structure leveraging the parity-check constraints.
+Let $\textbf{{A}}_{l}$ be a matrix in $\mathbb{R}^{N\times d_{l}}$, where $N$ is defined as $2n-k$ and $d_{l}$ as $n-k$.
+We then construct the extended parity-check matrix $\overline{\textbf{H}}$ as follows:
+$\overline{\textbf{H}}=\begin{bmatrix}{\textbf{H}}^{T}\\ {\textbf{{I}}}_{n-k}\end{bmatrix}$ (13)
+where ${\textbf{{I}}}_{n-k}\in\mathbb{R}^{(n-k)\times(n-k)}$ is an identity matrix with ones on the diagonal and zeros elsewhere.
+The generation of the attention mask is thus a direct process, formulated as $\text{M}(\overline{\textbf{H}})$.
+We naturally proceed to define the density of $\overline{\textbf{H}}$ as:
+$H_{d}=\frac{\text{sum}(\overline{\textbf{H}})}{(2n-k)\times(n-k)}$ (14)
 Figure 7: We present the corresponding matrices $\textbf{H}^{T}$, $\overline{\textbf{H}}$ and the Tanner graph for the Hamming (7,4) code.
 For enhanced clarity, Figure 7 offers a visual depiction of matrices $\textbf{H}^{T}$, $\overline{\textbf{H}}$ and the Tanner graph, utilizing the Hamming (7,4) code as an illustrative example.
-Here, matrix H is defined as
+Here, matrix H is defined as:
 $\textbf{H}=\begin{bmatrix}1&1&1&0&1&0&0\\ 1&0&1&1&0&1&0\\ 0&1&1&1&0&0&1\end{bmatrix}$ (15)
 Algorithm 2 Pseudocode for Sparse Mask Construction
-1 Input: H  # $shape=(n-k,n)$
-2 $H_{d}=\frac{sum(\textbf{H})}{(n-k)\times n}$
-3 $\overline{\textbf{H}}=zeros(2n-k,n-k)$
-4 $\overline{\textbf{H}}[0:n,0:n-k]=\textbf{H}.transpose(0,1)$
-5 $\overline{\textbf{H}}[n:2n-k,0:n-k]=rand(n-k,n-k)<H_{d}$
+1 Input: H  # $\text{shape}=(n-k,n)$
+2 $\overline{\textbf{H}}=\text{zeros}(2n-k,n-k)$
+3 $\overline{\textbf{H}}[0:n,0:n-k]=\textbf{H}.\text{transpose}(0,1)$
+4 $\overline{\textbf{H}}[n:2n-k,0:n-k]=\text{eye}(n-k)$
 5 Output: ()(¬$(-\infty)\cdot(\neg\mathrm{~{}\overline{\textbf{H}}})$
 This illustration demonstrates that the application of the mask results in sparser attention, a phenomenon particularly pronounced in low-density parity-check codes.
-Most importantly, the complexity of the UAM is now reduced to the density of the code $O(H_{d}\times d_{k}\times d_{l}\times N)$.
-Our experimental results indicate that the use of this mask has successfully reduced the computational complexity of the attention mechanism by an average of $83\%$.
+Most importantly, the complexity of the UAM is now reduced to the density of the code $O(N\times d_{l}\times d_{k}\times H_{d})$.
+Our experimental results indicate that the use of this mask has successfully reduced the computational complexity of the attention mechanism by an average of 86%.
 Furthermore, we provide a summary of the Python-style sparse mask construction within Algorithm 2.
-3.3 Architecture and Training
+3.3 Architecture and Training Methodology
 The architecture of the proposed unified error correction code Transformer is illustrated in Figure 8.
-The primary innovation in our model involves substituting the vanilla attention mechanism with a multi-head unified-attention module.
+The primary innovation in our model involves substituting the vanilla attention mechanism with a multi-head shared unified-attention module.
 This modification not only decreases computational complexity but also empowers the model to discern distinctive features among diverse codewords, thereby facilitating the establishment of a unified decoding architecture.
-We incorporate advanced pre-training strategies to optimize the model’s error correction performance.
+We employ fine-tuning techniques to enhance the model’s generalization to unseen codewords.
 Furthermore, we have designed standardization units to standardize the lengths of codewords and syndromes across diverse datasets, ensuring the model’s compatibility with various code types, lengths, and rates.
@@
 We employ binary cross-entropy as the loss function with the objective of learning to predict the multiplicative noise z.
-The corresponding target binary multiplicative noise is denoted as $=bin(sign(\textbf{{z}}))$ z ) ).
-Therefore, the loss calculation for a received individual codeword y is given by
-$loss=-\sum_{i=1}^{n}{\tilde{z}}_{i}log(f_{\theta}(y))+(1-{\tilde{z}}_{i})log(1-f_{\theta}(y))$ (17)
+The corresponding target binary multiplicative noise is denoted as z~=bin(sign($\textbf{{z}}))$ = bin ( sign ( z ) ).
+Therefore, the loss calculation for a received individual codeword y is given by:
+$\text{loss}=-\sum_{i=1}^{n}{\tilde{z}}_{i}\text{log}(f_{\theta}(y))+(1-{\tilde{z}}_{i})\text{log}(1-f_{\theta}(y))$ (16)
 […]
 $\widetilde{\textbf{{y}}}=\left[~{}\left[\lvert\textbf{{y}}\rvert,~{}\textbf{0}_{c}\right],~{}\left[s(\textbf{{y}}),~{}\textbf{0}_{s}\right]~{}\right]$ (17)
-where $\textbf{0}_{c}$ and $\textbf{0}_{s}$ represent zero vectors with sizes $N_{max}-size(\textbf{{y}})$ and $S_{max}-size(s(\textbf{{y}}))$, respectively.
+where $\textbf{0}_{c}$ and $\textbf{0}_{s}$ represent zero vectors with sizes $N_{\text{max}}-\text{size}(\textbf{{y}})$ and $S_{\text{max}}-\text{size}(s(\textbf{{y}}))$, respectively.
 Through this unit, the Transformer neural decoder can accommodate any codeword within the maximum length, allowing for mixed training in a unified manner.
-In the training phase, we conducted 1000 epochs, each consisting of 1024 minibatches, which in turn contained 512 samples each.
+In the training phase, we conducted 1000 epochs, each consisting of 1000 minibatches, which in turn contained 512 samples each.
 Samples for each minibatch were randomly selected from a pool of systematic Polar[31], LDPC, and BCH codewords.
@@
 We employ the Adam optimizer to dynamically adjust the learning rate, leveraging its efficacy in managing diverse data types and model architectures [33].
-The AWGN noise introduced to the links had its signal-to-noise ratio (SNR) randomly chosen between 4 and 6 for each batch.
-In the pre-training phase, we conducted 1000 epochs while maintaining consistency with all other parameters from the training phase.
-The word embeddings acquired during the pre-training stage served as the initial values for the subsequent training phase.
-The training and experiments were performed on NVIDIA GeForce RTX 4090 24GB GPUs.
+The AWGN noise introduced to the links had its Signal-to-Noise Ratio (SNR) randomly chosen between 3 and 7 for each batch.
+All experiments were conducted on NVIDIA GeForce RTX 4090 24GB GPUs.
 4 Experiments
@@
 The parity-check matrices utilized in the paper were sourced from [34].
-We benchmark our approach against the most recent state-of-the-art performance reported in [22], where each codeword parameter was trained individually and also contrast it with the legacy successive cancellation list decoding algorithm for Polar codes [32], as well as the normalized min-sum decoding algorithm for LDPC codes [10] [33].
+We benchmark our approach against the most recent state-of-the-art performance from [24], which trains each codeword parameter independently, and the hybrid-trained FECCT [25].
+Additionally, we compare it to legacy BP-based decoders, including unlearned BP [35], fully supervised Hyper BP [36], and Autoregressive BP (AR BP) [37].
 The results are presented in the form of bit error rate (BER) and block error rate (BLER) at various normalized SNR (i.e.
-$E_{b}/N_{0}$) values, with at least $10^{6}$ random codewords tested for each.
-The terms CBER and CBLER are hereby defined to denote the comprehensive bit error rate and block error rate, respectively, for the entirety of codewords utilized, spanning Polar, LDPC, and BCH codes.
-The proposed unified error correction code Transformer is defined by the following hyperparameters: the number of Transformer encoder layers $L$, the dimensions of word embeddings $d_{k}$, and the dimensions $d_{l}$ of the trainable memory $\textbf{{A}}_{l}$ and $\textbf{{V}}_{l}$ along with the number of unified attention heads $H$.
-We present the performance of our framework for four different architectures with $L\in\{3,6\},d_{k}\in\{16,32\}$, while all other hyperparameters are set to their default values: $d_{l}=64,H=8$.
-The hyperparameters in [22] are set as follows: $L=6,d_{k}=32,H=8$.
-Additionally, for Polar SCL decoding, the list size is set to 8, and for LDPC NMS decoding, the maximum number of iterations is set to 15 with a normalized factor of 0.875.
-Table 1 presents the simulation results, specifically showing the negative logarithm base 10 of the BER at $E_{b}/N_{0}$ values of 4, 5, and 6 dB.
-It is noted that higher values in this context are indicative of better performance.
-The best performance for each SNR is highlighted in bold and italics.
-We provide additional BER plots for Polar and LDPC codes in Figure 8 and Figure 9, respectively.
-Experimental results indicate that our approach achieves comparable performance to models trained individually as reported in [22], and outperforms both the SCL algorithm for Polar codes and the NMS algorithm for LDPC codes.
-Our proposed unified code-agnostic Transformer effectively decodes a wide range of code types, lengths, and rates.
-Especially for short LDPC codes, traditional NMS algorithms are constrained by a smaller girth, leading to poor performance because LLR information cannot propagate effectively during the iteration process.
-In contrast, our approach sidesteps this issue by eliminating iteration, thereby offering a practical solution for short LDPC code applications, such as those for future 6G mobile communication control channel coding schemes.
-Furthermore, the performance gains observed with smaller dimensions $d_{l}=64\ll N=192$ confirm that the proposed decoding architecture substantially reduces computational complexity.
-Table 1: A comparison of the negative logarithm base 10 of the BER
-Figure 8: BER for Polar codes under $L=6,d_{k}=32$.
-Figure 9: BER for LDPC codes under $L=6,d_{k}=32$.
+$E_{b}/N_{0}$) values, with at least $10^{5}$ random codewords tested per code type.
+We define aggregate BER (ABER) and aggregate BLER (ABLER) as the overall error rates spanning all evaluated codewords.
+In addition, to enable direct comparison with ECCT and FECCT, whose simulation results were directly obtained from their respective papers [24, 25], we fine-tuned the pre-trained UECCT models on Consultative Committee for Space Data Systems (CCSDS) [38] and MacKay [39] codewords leveraging the hybrid-trained Polar, LDPC, and BCH models.
+Given GPU memory constraints and the fact that the performance of long codes has been thoroughly explored [40], while the error correction of short codes remains far from the Shannon limit [41], this work primarily focuses on training and validating medium-to-short codes.
+The proposed unified error correction code Transformer is defined by the following hyperparameters: the number of Transformer encoder layers $L$, the number of unified attention heads $H$, the dimensions of word embeddings $d_{k}$, the dimensions $d_{l}$ of the trainable memory $\textbf{{A}}_{l}$ and $\textbf{{V}}_{l}$, and the dimension $d_{f}$ of the FFN layer.
+We present the performance of our framework with hyperparameters are set to $L=6$, $H=8$, $d_{k}=64$, $d_{l}=64$, and $d_{f}=4\cdot H\cdot d_{k}$.
+The value of $d_{l}=64$ corresponds to the maximum syndrome length among the trained codewords.
+Figure 9: BER comparison of ECCT, FECCT, and proposed UECCT for Polar codes.
+Figure 10: BER comparison of ECCT, FECCT, and proposed UECCT for BCH codes.
+Table 1: A comparison of the negative natural logarithm of the BER
+Table 1 presents the simulation results, specifically showing the negative natural logarithm of the BER at $E_{b}/N_{0}$ values of 4, 5, and 6 dB.
+It is noted that higher values in this context indicate better performance, with the best results for each SNR highlighted in bold.
+Additional BER plots for Polar and BCH codes are provided in Figure 9 and Figure 10, respectively.
+Simulation results demonstrate that our proposed UECCT outperforms both the individually trained ECCT [24] and the hybrid-trained FECCT [25] across the majority of codewords, validating the effectiveness of the proposed architecture.
+Notably, UECCT even surpasses FECCT’s performance with an embedding length of 128 when operating at a more compact embedding length of 64, further highlighting the efficiency and power of our design.
+The superior performance stems from its ability to leverage a shared attention mechanism across heads, as detailed in Section 3, combined with sparsity-aware masking that reduces computational complexity while preserving decoding accuracy.
 5 Analysis
-We delve into the influence of Transformer encoder layers on overall performance, examining both the role of the dimension parameter $d_{k}$ and the efficacy of sparse masked attention mechanisms.
-Furthermore, we conduct a comparative analysis of the computational complexity inherent in our methodology versus that of alternative approaches.
-5.1 Impact of the Transformer Encoder Layers and the Dimensions $d_{k}$
-Figure 10 illustrates the CBER performance across various Transformer encoder layers and the dimensions $d_{k}$.
-It can be observed that the number of Transformer encoder layers exerts a greater impact on performance than the dimension $d_{k}$.
-This increase in depth effectively enlarges the parameter search space, potentially enhancing the model’s nonlinear fitting capability and enabling more complex expression.
-The results suggest that a deeper Transformer architecture is more capable of capturing intricate patterns and dependencies within the data, which is crucial for improving error correction performance.
-By adding more layers, the model can perform multiple levels of abstraction and refinement, allowing it to better understand and correct errors.
-This is especially crucial in the context of wireless communication systems, which are characterized by complexity and variability that simple models might not adequately capture.
-In essence, the findings emphasize the importance of model depth in Transformer architectures for error correction tasks.
-This insight can guide the design of more effective models by prioritizing the depth of the encoder layers over the dimension of $d_{k}$, leading to better performance in practical applications.
-Figure 10: CBER performance across various Transformer encoder layers and the dimensions $d_{k}$.
-5.2 Impact of the Sparse Masked Attention
-We illustrate the impact of sparse masked attention on loss, CBER, and CBLER in Fig.
-11.
-Specifically, sparse masked attention reduces the loss by $45\%$ and leads to faster convergence during training.
-Furthermore, this approach significantly improves error correction performance: the CBER shows an improvement of approximately $0.4$ dB at a bit error rate of $10^{-5}$, and the CBLER improves by about $0.5$ dB at a block error rate of $10^{-4}$.
-These improvements underscore the efficacy of integrating domain-specific knowledge into the attention mechanism, demonstrating substantial gains in both efficiency and accuracy.
-This approach not only reduces computational complexity but also enhances the model’s ability to decode complex signals, making it a promising strategy for future advancements in error correction techniques.
-Figure 11: The overall improvement in loss, CBER and CBLER with sparse masked attention.
+We investigate the impact of Transformer encoder layers on performance, focusing on the embedding dimension $d_{k}$ and the effectiveness of sparse masked attention.
+Additionally, we compare the computational complexity of our approach with existing methods.
+5.1 Impact of Encoder Layers and Embedding Dimensions
+Figure 11 displays the ABER performance across varying encoder layers $L$ and embedding dimensions $d_{k}$.
+Results reveal that $L$ exerts a stronger influence on performance than $d_{k}$.
+Increasing $L$ broadens the parameter space, enhancing nonlinear fitting and enabling complex representations.
+Figure 11: ABER performance across various Transformer encoder layers and the embedding dimensions $d_{k}$.
+Deeper architectures better capture intricate data patterns and dependencies, crucial for error correction.
+Additional layers enable multi-level abstraction and refinement, improving error interpretation and correction, especially in complex wireless communication systems where simpler models falter.
+These findings highlight the primacy of model depth in Transformer-based error correction. Prioritizing $L$ over $d_{k}$ can optimize model design, enhancing practical performance.
+5.2 Impact of Sparse Masked Attention
+Figure 12 illustrates the effect of sparse masked attention on training loss, ABER, and ABLER.
+Sparse masked attention reduces training loss by 33.3% and speeds up convergence.
+It also significantly boosts error correction, improving ABER by 0.7 dB at a bit error rate of $10^{-5}$ and ABLER by 0.9 dB at a block error rate of $10^{-4}$.
+Figure 12: The overall improvement in loss, ABER and ABLER with sparse masked attention.
+These gains highlight the effectiveness of embedding domain-specific knowledge into the attention mechanism, yielding significant improvements in efficiency and accuracy.
+By reducing computational complexity while boosting the model’s capacity to decode complex signals, this approach emerges as a promising advancement for future error correction techniques.
 5.3 Computational Complexity Analysis
@@
 The principal computational complexity within each layer is found in the multi-head self-attention module, which is predominantly responsible for matrix multiplication and is defined by the hyperparameters $H$, $d_{k}$ and $d_{l}$.
-Consequently, the computational complexity of the proposed architecture is given by $O\left({L\times H\times\left({{H_{d}}\times{d_{k}}\times{d_{l}}\times N+d_{k}^{2}}\right)}\right)$, where $N=2n-k$, $d_{l}\ll N$ and $H\ll N$.
+Consequently, the computational complexity of the proposed architecture is given by $O\left(L\times H\times\left(N\times d_{l}\times d_{k}\times H_{d}\right)\right)$, where $N=2n-k$, and $d_{l}\ll N$.
 Thanks to the sparsity of the parity-check matrix, $H_{d}$ is usually quite small.
-In our experiment, $H_{d}$ is approximately 0.17, which significantly reduces the computational complexity of the neural network.
-In comparison, the approach in [22], which employs the vanilla Transformer architecture, has a complexity of $O\left({L\times H\times\left({{N^{2}}\times{d_{k}}+d_{k}^{2}}\right)}\right)$.
-When it comes to non-learning solutions, the computational complexities of the SCL and NMS decoding algorithms are $O\left(L_{s}\times n\times log_{2}\left(n\right)\right)+O\left(L_{s}\times\left(n-1\right)\right)+k\times O\left(2\times L_{s}\times log_{2}\left(2\times L_{s}\right)\right)$ and $I_{max}\times O\left(2\times n\times d_{\nu}+M\times\left(2\times d_{c}-1\right)\right)+I_{max}\times O\left(M\times d_{c}\right)$, respectively [34].
-Here, $L_{s}$ is the list size, $I_{max}$ is the maximum number of iterations, $M$ is the size of the check matrix, $d_{v}$ is the number of variable node, and $d_{c}$ is the number of parity check node.
-Compared to the approach in [22], while the proposed method excels in decoding versatility and computational complexity, it may fall short in terms of memory efficiency, power consumption, and computational resource utilization when compared to non-learning solutions.
+In our experiment, $H_{d}$ is approximately 0.14, which significantly reduces the computational complexity of the neural network.
+By contrast, the vanilla Transformer architectures in ECCT and FECCT have complexities of $O\left(L\times H\times\left(N^{2}\times d_{k}\times M_{d}+N\times d_{k}^{2}\right)\right)$ and $O\left(L\times H\times\left(N^{2}\times d_{k}+N\times d_{k}^{2}+N^{2}\times M_{d}\right)\right)$, respectively, where $M_{d}$ represents the mask density in these models.
+Table 2: Comparison of trainable parameters, MACs, training time, inference time, and mask density among ECCT, FECCT, and UECCT for various codes, all with identical parameters $L=6$, $H=8$, $d_{k}=64$, $d_{l}=64$, and $d_{f}=2048$
+Table 2 compares the Trainable Parameters (TPs), Multiply-Accumulate Operations (MACs), training time, inference time, and mask density for ECCT, FECCT, and UECCT.
+Lower values are preferable, with the best performing values highlighted in bold.
+The metrics for TPs and MACs are obtained using Thop [42], a neural network profiling tool.
+UECCT reduces parameters and MACs by 25.0% compared to both ECCT and FECCT through low-rank approximation, shared attention, and omission of Q and K projections.
+This results in training times 19.7% and 81.2% lower, and inference times 23.1% and 93.5% lower than ECCT and FECCT, respectively.
+Mask densities are influenced by design choices: FECCT has the highest density due to Tanner graph distance calculations, ECCT has a moderate density from extended bit-pair relations, and UECCT has the lowest density, 58.2% and 79.7% lower than ECCT and FECCT, achieved through the direct utilization of H and the extension of the identity matrix.
+For a more intuitive representation, the average values of each metric across the codewords in Table 2 are provided in Figure 13.
+Figure 13: Comparison of average TPs, MACs, training time, inference time, and mask density between ECCT, FECCT and UECCT.
+Compared to the approach in [24, 25], while the proposed method excels in decoding versatility and computational complexity, it may fall short in terms of memory efficiency, power consumption, and computational resource utilization when compared to non-learning solutions.
 This inefficiency could limit its potential for deployment.
@@
 6 Conclusions
-We introduced a novel, unified code-agnostic Transformer decoding architecture, designed to integrate various linear block codes seamlessly and offer superior decoding performance.
+We introduced a unified, code-agnostic Transformer decoding architecture, designed to integrate various linear block codes seamlessly and offer superior decoding performance.
 Our approach leverages the self-attention mechanism inherent in Transformer models, enabling effective handling of sequential data and capturing long-range dependencies critical for decoding tasks.

### v3

@@
     High Computational Efficiency: The unified attention module in our proposed Transformer-based decoding architecture demonstrates linear computational complexity by eliminating the need to project the input into higher dimensions.
-This results in a significant reduction in complexity from the traditional $O(N^{2})$ to a more efficient $O(N)$, where $N$ represents the sequence length.
+This results in a significant reduction in complexity from the traditional $\mathcal{O}(N^{2})$ to a more efficient $\mathcal{O}(N)$, where $N$ represents the sequence length.
 Additionally, by employing our sparse mask, we further achieve an $86\%$ reduction in computational complexity.
@@
 $\mathbb{F}_{2}$ represents the Galois field of order 2, and $\mathbb{R}$, $\mathbb{N}$ denote the real and natural numbers.
-The transpose is indicated by $[\cdot]^{T}$, and $\mathcal{N}(\mu,\sigma^{2})$ denotes a Gaussian distribution with mean $\mu$ and variance $\sigma^{2}$.
-The function $\text{bipolar}(x)=1-2x$ maps $x\in{0,1}$ to bipolar values ${1,-1}$, with its inverse $\text{bin}(x)=\frac{1-x}{2}$ mapping ${1,-1}$ back to ${0,1}$.
-The operator $\text{size}(\cdot)$ denotes the number of elements in a vector or matrix, and $\text{sign}(\cdot)$ represents the sign function, defined as $\text{sign}(x)=1$ if $x>0$, 00 if $x=0$, and $-1$ if $x<0$.
+The transpose is indicated by $\cdot]^{T}$, and $\mathcal{N}(\mu,\sigma^{2})$ denotes a Gaussian distribution with mean $\mu$ and variance $\sigma^{2}$.
+The function $\text{bipolar}(x)=1-2x$ maps $x\in{0,1}$ to bipolar values 1,1{1,-1}, with its inverse $\text{bin}(x)=\frac{1-x}{2}$ mapping 1,1{1,-1} back to 0,1{0,1}.
+The operator $\text{size}(\cdot)$ denotes the number of elements in a vector or matrix, and $\text{sign}(\cdot)$ represents the sign function, defined as $\text{sign}(x)=1$ if $x>0$, 0 if $x=0$, and 1-1 if $x<0$.
 For a linear block code with code length $n$ and information bit length $k$, we denote it as $(n,k)$.
@@
 $\text{MHA}\left(\textbf{{Q}},\textbf{{K}},\textbf{{V}}\right)=\text{Concat}\left(\text{head}_{1},\cdots,\text{head}_{H}\right)\cdot\textbf{{W}}^{\textit{O}}$ (3)
-where $\text{head}_{i}=\text{Attn}(\textbf{{QW}}_{i}^{Q},\textbf{{KW}}_{i}^{K},\textbf{{VW}}_{i}^{V})$, with $\textbf{{Q}},\textbf{{K}},\textbf{{V}}\in\mathbb{R}^{N\times d_{k}}$ set to $\textbf{{X}}_{e}$ for self-attention, using trainable weights $\textbf{{W}}_{i}^{Q},\textbf{{W}}_{i}^{K},\textbf{{W}}_{i}^{V}\in\mathbb{R}^{d_{k}\times d_{k}}$ and $\textbf{{W}}^{O}\in\mathbb{R}^{Hd_{k}\times Hd_{k}}$, as depicted in Figure 2.
+where $\text{head}_{i}=\text{Attn}(\textbf{{X}}_{Q}\textbf{{W}}_{i}^{Q},\textbf{{X}}_{K}\textbf{{W}}_{i}^{K},\textbf{{X}}_{V}\textbf{{W}}_{i}^{V})$, with $\textbf{{X}}_{Q},\textbf{{X}}_{K},\textbf{{X}}_{V}\in\mathbb{R}^{N\times d_{k}}$ set to $\textbf{{X}}_{e}$ for self-attention, using trainable weights $\textbf{{W}}_{i}^{Q},\textbf{{W}}_{i}^{K},\textbf{{W}}_{i}^{V}\in\mathbb{R}^{d_{k}\times d_{k}}$ and $\textbf{{W}}^{O}\in\mathbb{R}^{Hd_{k}\times Hd_{k}}$, as depicted in Figure 2.
 Figure 2: Architecture of multi-head self-attention.
@@
 where $\textbf{{W}}_{1}\in\mathbb{R}^{d_{f}\times d_{h}}$, $\textbf{{W}}_{2}\in\mathbb{R}^{d_{h}\times d_{f}}$, $\textbf{{b}}_{1}$, $\textbf{{b}}_{2}$ are trainable parameters, $d_{h}=H\cdot d_{k}$ and $\text{ReLU}(x)=\max(0,x)$ enables complex feature learning, with $d_{f}>d_{h}$ for enhanced performance.
-Finally, residual connections and layer normalization link the MHA and FFN layers:
+Finally, residual connections and layer normalization link the MHA and FFN layers to produce the output $\textbf{{X}}_{2}$ of a single-layer Transformer encoder:
 $\textbf{{X}}_{1}$ $=\text{MHA}\left(\text{LayerNorm}\left(\textbf{{X}}_{e}\right)\right)+\textbf{{X}}_{e}$ (5)
@@
 We adopt a linear block code defined by a generator matrix $\textbf{G}\in\mathbb{R}^{k\times n}$ and a parity-check matrix $\textbf{H}\in\mathbb{R}^{(n-k)\times n}$, satisfying $\textbf{G}\cdot\textbf{H}^{T}=\textbf{0}$ over $\mathbb{F}_{2}$.
-[…]
+The encoder maps a message vector $\textbf{{m}}\in\{0,1\}^{k}$ to a codeword $\textbf{{x}}\in\{0,1\}^{n}$ via $\textbf{{x}}=\textbf{{m}}\cdot\textbf{G}$, where $\textbf{H}\cdot\textbf{{x}}=\textbf{0}$.
+This codeword is modulated using Binary Phase Shift Keying (BPSK), defined as $\text{BPSK}(x)=1-2x$, yielding $\textbf{{x}}_{s}$, which is transmitted over Binary-Input Discrete Memoryless Channels (B-DMCs, e.g., AWGN), producing $\textbf{{y}}=\textbf{{x}}_{s}+\textbf{{n}}$, with $\textbf{{n}}\sim\mathcal{N}(0,\sigma^{2})$.
 Figure 3: Architecture of the communication system.
@@
 In the pre-processing step, the absolute values of received signal and the syndromes are concatenated as the input to the neural network, which can be written as:
-y~=[|y|,s($\textbf{{y}}\right)\right]$ = [ | y | , s ( y ) ] (7)
+$\widetilde{\textbf{{y}}}=\left[\lvert\textbf{{y}}\rvert,\text{s}\left(\textbf{{y}}\right)\right]$ (7)
+where, $\left[\cdot,\cdot\right]$ denotes the concatenation of vectors, $\lvert\textbf{{y}}\rvert$ signifies the absolute value of y and $\text{s}(\textbf{{y}})=\textbf{H}\cdot\text{bin}(\text{sign}(\textbf{{y}}))\in\{0,1\}^{n-k}$ is the syndrome from hard-decoded y.
+In the post-processing step, the predicted noise $\hat{\textbf{{z}}}$ is multiplied by the received signal y to recover x. That is, the predicted $\hat{\textbf{{x}}}$ takes the form:
 […]
-In the post-processing step, the predicted noise $\hat{\textbf{{z}}}$ is multiplied by the received signal y to recover x. That is, the predicted $\hat{\textbf{{x}}}$ takes the form:
-$\hat{\textbf{{x}}}=\text{bin}\left(\text{sign}\left({\textbf{{y}}}\cdot{\hat{\textbf{{z}}}}\right)\right)$ (8)
 2.3 Mask Generation in ECCT and FECCT
@@
 3.1 Shared Unified Attention
-The vanilla self-attention module can be viewed as establishing correlations among different input features within a data sample while neglecting the intrinsic relationships between different samples.
-This limitation hinders its ability to learn the distinctiveness of features between various code types in channel decoding tasks.
-Furthermore, research in [30] suggests that the input of the softmax function is probabilistically sparse, indicating that only a few weights dominate.
-This implies that an $N\times N$ self-attention matrix is not strictly necessary.
-This sparsity suggests that a low-rank approximation could effectively capture the dominant correlations without sacrificing performance.
+The vanilla self-attention mechanism in Transformer architectures establishes dense correlations among all input features within a sequence, resulting in a computational complexity of $\mathcal{O}(N^{2}\cdot d_{k}+N\cdot d_{k}^{2})$, where $N$ is the sequence length and $d_{k}$ is the embedding dimension.
+This approach, while effective for capturing long-range dependencies, is computationally intensive and overlooks the intrinsic relationships across different samples, particularly in channel decoding tasks involving diverse linear block codes.
+Moreover, prior studies [30] suggest that the input to the softmax function in self-attention is probabilistically sparse, with only a few weights dominating the attention distribution.
+This sparsity implies that a low-rank approximation could effectively capture the dominant correlations, reducing redundancy without compromising performance.
 Figure 5: Rank distribution of the $\textbf{{Q}}\cdot\textbf{{K}}^{T}$ matrix in the ECCT model across 6 encoder layers for various code types.
-To validate this, we trained an ECCT model (with parameters $L=6$, $H=8$, $d_{k}=32$, $d_{f}=1024$) on a variety of linear block codes, including Polar(32, 16), Polar(64, 32), Polar(128, 86), LDPC(49, 24), LDPC(121, 60), LDPC(121, 80), BCH(31, 16), BCH(63, 36), and BCH(127, 120).
-We analyzed the rank distribution of the $\textbf{{Q}}\cdot\textbf{{K}}^{T}$ matrix across 1000 batches for the 6-layer Transformer encoder.
-As shown in Figure 5, the rank ranges from 11 to 32, consistently indicating a low-rank structure across all layers and code types.
-Based on the aforementioned considerations, we have developed a novel, low-complexity Unified Attention Module (UAM) to tackle channel decoding tasks involving multiple code types, lengths, and rates.
-We introduced learnable memory components, $\textbf{{A}}_{l}$ (attention) and $\textbf{{V}}_{l}$ (values), which play the role of learning features from the entire training dataset.
-The mathematical expression is given by:
+To investigate this hypothesis, we trained an ECCT model [24] with parameters $L=6$ (number of Transformer encoder layers), $H=8$, $d_{k}=\text{max}(2n-k)=182$, and $d_{f}=2912$ on a diverse set of linear block codes, including Polar(32,16), Polar(64,32), Polar(128,86), LDPC(49,24), LDPC(121,60), LDPC(121,80), BCH(31,16), BCH(63,36), and BCH(127,120).
+We analyzed the rank distribution of the $\textbf{{Q}}\cdot\textbf{{K}}^{T}$ matrix across 1000 batches for the six-layer Transformer encoder.
+As illustrated in Figure 5, the rank ranges from 12 to 181, with over 90% of values below 69, consistently indicating a low-rank structure across all layers and code types.
+This observation suggests that the high-dimensional self-attention matrix can be approximated with a lower-rank representation, offering a pathway to reduce computational overhead.
+Figure 6: Distribution of Jensen-Shannon divergence between attention head pairs across 6 layers, showing high similarity (JSD << 0.17).
+Furthermore, the conventional multi-head self-attention mechanism computes independent attention scores for each head using distinct weight matrices $\textbf{{W}}_{i}^{Q}$, $\textbf{{W}}_{i}^{K}$, and $\textbf{{W}}_{i}^{V}$.
+However, we hypothesize that the attention distributions across heads may exhibit high similarity, particularly in channel decoding, where the sparse structure of the parity-check matrix $\mathbf{H}$ constrains bit interactions.
+To quantify this, we employed the Jensen-Shannon Divergence (JSD) [31], a symmetric measure of similarity between probability distributions, defined as:
+$\text{JSD}(\mathbf{A}_{i}[k]\|\mathbf{A}_{j}[k])=\frac{1}{2}\left[D_{KL}(\mathbf{A}_{i}[k]\|\mathbf{M}[k])+D_{KL}(\mathbf{A}_{j}[k]\|\mathbf{M}[k])\right]$ (9)
+where $k$ is the row index, $\mathbf{A}_{i}$ and $\mathbf{A}_{j}$ $\in\mathbb{R}^{N\times N}$are the softmax outputs of attention heads $i$ and $j$, $\mathbf{M}[k]=\frac{1}{2}(\mathbf{A}_{i}[k]+\mathbf{A}_{j}[k])$ is the average distribution, and $D_{KL}(\mathbf{P}\|\mathbf{Q})=\sum_{l=1}^{N}\mathbf{P}[l]\log\frac{\mathbf{P}[l]}{\mathbf{Q}[l]}$ is the Kullback-Leibler divergence [32].
+Using the same simulation setup as the low-rank validation ($L=6$, $H=8$, $d_{k}=182$, $d_{f}=2912$), we computed the per-row JSD for all head pairs across all rows, averaging the results.
+As shown in Figure 6, JSD values are consistently below 0.17 across all layers, with a mean of 0.13, indicating strong alignment in attention distributions.
+This similarity, driven by the sparse structure of $\mathbf{H}$, supports the feasibility of sharing a common attention matrix across heads.
+Motivated by these findings, we propose a novel Unified Attention Module (UAM) to address the computational inefficiency and lack of cross-code generalization in traditional self-attention.
+The UAM introduces learnable memory components, $\textbf{{A}}_{l}$ and $\textbf{{V}}_{l}\in\mathbb{R}^{N\times d_{l}}$, where $d_{l}$ is a hyperparameter controlling the low-rank approximation.
+The UAM is defined as:
 $\text{UniAttn}\left({{\textbf{{A}}_{l}},{\textbf{{V}}_{l}}}\right)=\text{Softmax}\left({{\textbf{{A}}_{l}}}\right)\cdot\left({{\textbf{{V}}_{l}^{T}}\textbf{{X}}}\right)$ (10)
-where $\textbf{{X}}\in\mathbb{R}^{N\times d_{k}}$ is the input of attention layer, $\textbf{{A}}_{l},\textbf{{V}}_{l}\in\mathbb{R}^{N\times d_{l}}$ are trainable parameters, and $d_{l}$ is an adjustable hyperparameter controlling the low-rank approximation.
-In order to gain a better understanding of the proposed UAM, we rewrite equation (2) as follows:
+where $\textbf{{X}}\in\mathbb{R}^{N\times d_{k}}$ is the input to the attention layer. In contrast, the conventional self-attention is expressed as:
 $\text{Attn}\left(\textbf{{W}}^{\textit{Q}},\textbf{{W}}^{\textit{K}},\textbf{{W}}^{\textit{V}}\right)=\text{Softmax}\left(\frac{\left(\textbf{{XW}}^{\textit{Q}}\right)\cdot{\left(\textbf{{XW}}^{\textit{K}}\right)}^{T}}{\sqrt{d_{k}}}\right)\cdot\left(\textbf{{XW}}^{\textit{V}}\right)$ (11)
+By eliminating the projections $\textbf{{W}}^{\textit{Q}}$ and $\textbf{{W}}^{\textit{K}}$, the UAM reduces the computational complexity from $\mathcal{O}(N^{2}\cdot d_{k}+N\cdot d_{k}^{2})$ to $\mathcal{O}(N\cdot d_{l}\cdot d_{k})$, where $d_{l}\ll N$.
+The shared attention matrix $\mathbf{A}_{l}$ is applied across all heads, leveraging the observed head similarity to foster a cohesive decoding architecture.
+This design not only reduces the parameter count but also enables the model to learn shared features across diverse code types, facilitating a code-agnostic framework.
 Figure 7: Architecture of the vanilla multi-head self-attention (A) and the proposed multi-head unified-attention (B).
@@
 6 $\textbf{{A}}=\text{Softmax}(\textbf{{A}}_{l},\text{dim}=-1)$  # $\text{shape}=(B,1,N,d_{l})$
-7 $_{o}=\textbf{{A}}\cdot\left(\textbf{{V}}_{l}^{T}\cdot\textbf{{X}}\right)$ X )  # $\text{shape}=(B,H,N,d_{k})$
+7 $\textbf{{X}}_{o}=\textbf{{A}}\cdot\left(\textbf{{V}}_{l}^{T}\cdot\textbf{{X}}\right)$  # $\text{shape}=(B,H,N,d_{k})$
 8 $\textbf{{X}}_{o}=\textbf{{X}}_{o}.\text{transpose}(1,2).\text{view}(B,N,d_{h})$
@@
 10 Output: $\textbf{{X}}_{o}$  # $\text{shape}=(B,N,d_{h})$
-By comparing equations (9) and (10), the computational complexity of the vanilla self-attention module is $O(N^{2}\times d_{k}+N\times d_{k}^{2})$, whereas the proposed UAM reduces it to $O(N\times d_{l}\times d_{k})$.
-Importantly, the omission of the Q and K projections in the vanilla Transformer leads to significant parameter savings.
-As demonstrated in subsequent experiments, a $d_{l}$ value comparable to $n-k$ is sufficient to achieve a favorable bit error rate.
-This is supported by the fact that the maximum rank of a parity-check matrix $\textbf{H}\in\{0,1\}^{(n-k)\times n}$ is $n-k$, providing a theoretical basis for the observed performance.
-It is also important to note that both ECCT and FECCT are based on the vanilla Transformer architecture, differing only in their self-attention mask designs.
-The proposed multi-head unified-attention and corresponding Python-style pseudocode are illustrated in Figure 6 and Algorithm 1, respectively.
-As depicted in Figure 6, thanks to the sparsity of the attention scores, multiple heads are able to share the same attention $\textbf{{A}}_{l}$, resulting in a more cohesive decoding architecture and further reducing the complexity of the neural network.
-This approach contrasts with vanilla self-attention mechanisms, in which each head independently computes its scores using separate learned weights.
+The architecture of the proposed multi-head UAM is depicted in Figure 7, with the corresponding pseudocode provided in Algorithm 1.
+The shared attention mechanism, combined with the low-rank approximation, significantly reduces model complexity while preserving decoding accuracy, as validated in subsequent experiments.
+This approach marks a significant advancement over the vanilla Transformer-based ECCT [24], which lacks cross-code unification.
 3.2 Sparse Masked Unified Attention
@@
 $\text{UniAttn}\left(\boldsymbol{A}_{l},\boldsymbol{V}_{l}\right)=\text{Softmax}\left(\boldsymbol{A}_{l}+\text{M}(\textbf{H})\right)\cdot\left(\boldsymbol{V}_{l}^{T}\boldsymbol{X}\right)$ (12)
-[…]
+The design of the mask function $\text{M}(\textbf{H})$ aims to enable the neural network to capture the relationship between the received signal $\textbf{{y}}\in\mathbb{R}^{n}$ and its syndrome $\text{s}(\textbf{{y}})=\textbf{H}\cdot\overline{\textbf{{y}}}\in\{0,1\}^{n-k}$, where $\overline{\textbf{{y}}}=\text{bin}(\text{sign}(\textbf{{y}}))$ is the hard-decoded binary vector derived from y.
+For a valid codeword y (i.e., noise-free or correctly decoded), the orthogonality condition $\textbf{H}\cdot\overline{\textbf{{y}}}=\textbf{0}$ holds over $\mathbb{F}_{2}$, yielding $\text{s}(\textbf{{y}})=\textbf{0}$.
+To embed this relationship within the attention mechanism, we define an extended parity-check matrix $\overline{\textbf{H}}\in\mathbb{R}^{(2n-k)\times(n-k)}$ satisfying:
 $\left[\overline{\textbf{{y}}},\text{s}\left(\textbf{{y}}\right)\right]\cdot\overline{\textbf{H}}=\textbf{0}$ (13)
 Inspired by this, we explore a mask structure leveraging the parity-check constraints.
-Let $\textbf{{A}}_{l}$ be a matrix in $\mathbb{R}^{N\times d_{l}}$, where $N$ is defined as $2n-k$ and $d_{l}$ as $n-k$.
+Let $\textbf{{A}}_{l}$ be a matrix in $\mathbb{R}^{N\times d_{l}}$, where $N$ is defined as 2nk2n-k and $d_{l}$ as nkn-k.
+The theoretical foundation for setting $d_{l}\approx n-k$ stems from the maximum rank of the parity-check matrix $\textbf{H}\in\mathbb{F}_{2}^{(n-k)\times n}$, which is nkn-k.
 We then construct the extended parity-check matrix $\overline{\textbf{H}}$ as follows:
@@
 4 $\overline{\textbf{H}}[n:2n-k,0:n-k]=\text{eye}(n-k)$
-5 Output: ()(¬$(-\infty)\cdot(\neg\mathrm{~{}\overline{\textbf{H}}})$
+5 Output: $(-\infty)\cdot(\neg\mathrm{~\overline{\textbf{H}}})$
 This illustration demonstrates that the application of the mask results in sparser attention, a phenomenon particularly pronounced in low-density parity-check codes.
-Most importantly, the complexity of the UAM is now reduced to the density of the code $O(N\times d_{l}\times d_{k}\times H_{d})$.
+Most importantly, the complexity of the UAM is now reduced to the density of the code $\mathcal{O}(N\times d_{l}\times d_{k}\times H_{d})$.
 Our experimental results indicate that the use of this mask has successfully reduced the computational complexity of the attention mechanism by an average of 86%.
 Furthermore, we provide a summary of the Python-style sparse mask construction within Algorithm 2.
+It should be noted that, in Application-Specific Integrated Circuit (ASIC) implementations, computations involving the masked parts can be directly skipped to enhance efficiency.
 3.3 Architecture and Training Methodology
@@
 Figure 9: Illustration of the proposed unified error correction code Transformer.
-In Figure 8, the output y from the AWGN channel passes through the pre-processing module to yield $N=2n-k$ input elements for the neural network.
+In Figure 9, the output y from the AWGN channel passes through the pre-processing module to yield N=2nkN=2n-k input elements for the neural network.
 The decoder is constructed by concatenating $L$ Transformer encoders, which consist of unified-attention and feed-forward layers, with normalization layers in between.
 The output module is characterized by two fully connected layers.
-The initial layer compresses the multi-head output into a one-dimensional vector of size $2n-k$, with the subsequent layer transforming it into an $n$-dimensional vector representing the softly decoded noise.
+The initial layer compresses the multi-head output into a one-dimensional vector of size 2nk2n-k, with the subsequent layer transforming it into an $n$-dimensional vector representing the softly decoded noise.
 Finally, the post-processing module acts on the estimated noise to derive the transmitted original codeword.
 We employ binary cross-entropy as the loss function with the objective of learning to predict the multiplicative noise z.
-The corresponding target binary multiplicative noise is denoted as z~=bin(sign($\textbf{{z}}))$ = bin ( sign ( z ) ).
+The corresponding target binary multiplicative noise is denoted as $\tilde{\textbf{{z}}}=\text{bin}(\text{sign}(\textbf{{z}}))$.
 Therefore, the loss calculation for a received individual codeword y is given by:
 $\text{loss}=-\sum_{i=1}^{n}{\tilde{z}}_{i}\text{log}(f_{\theta}(y))+(1-{\tilde{z}}_{i})\text{log}(1-f_{\theta}(y))$ (17)
-[…]
+To accommodate codewords of varying lengths within a unified architecture, we have designed a standardization unit that performs padding on $\lvert\textbf{{y}}\rvert$ and $s(\textbf{{y}})$ as described in Equation (7).
+Specifically, after receiving the channel output, we pad different codewords and their corresponding syndromes with zeros up to the maximum length.
+Assuming that the maximum length of the codeword is $N_{\text{max}}$ and the maximum length of the syndrome is $S_{\text{max}}$, the dataset is standardized according to:
 $\widetilde{\textbf{{y}}}=\left[~\left[\lvert\textbf{{y}}\rvert,~\textbf{0}_{c}\right],~\left[s(\textbf{{y}}),~\textbf{0}_{s}\right]~\right]$ (18)
@@
 Through this unit, the Transformer neural decoder can accommodate any codeword within the maximum length, allowing for mixed training in a unified manner.
+In ASIC implementations, computations for zero-padded parts can be skipped to improve efficiency, similar to the sparse-masked attention.
 In the training phase, we conducted 1000 epochs, each consisting of 1000 minibatches, which in turn contained 512 samples each.
@@
 Figure 10: BER comparison of ECCT, FECCT, and proposed UECCT for Polar codes.
+Figure 11: BER comparison of ECCT and proposed UECCT for LDPC codes.
 Figure 12: BER comparison of ECCT, FECCT, and proposed UECCT for BCH codes.
@@
 It is noted that higher values in this context indicate better performance, with the best results for each SNR highlighted in bold.
-Additional BER plots for Polar and BCH codes are provided in Figure 9 and Figure 10, respectively.
+Additional BER plots for Polar, LDPC and BCH codes are provided in Figure 10, 11 and Figure 12, respectively.
 Simulation results demonstrate that our proposed UECCT outperforms both the individually trained ECCT  [24] and the hybrid-trained FECCT  [25] across the majority of codewords, validating the effectiveness of the proposed architecture.
@@
 The principal computational complexity within each layer is found in the multi-head self-attention module, which is predominantly responsible for matrix multiplication and is defined by the hyperparameters $H$, $d_{k}$ and $d_{l}$.
-Consequently, the computational complexity of the proposed architecture is given by $O\left(L\times H\times\left(N\times d_{l}\times d_{k}\times H_{d}\right)\right)$, where $N=2n-k$, and $d_{l}\ll N$.
+Consequently, the computational complexity of the proposed architecture is given by $\mathcal{O}\left(L\times H\times\left(N\times d_{l}\times d_{k}\times H_{d}\right)\right)$, where N=2nkN=2n-k, and $d_{l}\ll N$.
 Thanks to the sparsity of the parity-check matrix, $H_{d}$ is usually quite small.
 In our experiment, $H_{d}$ is approximately 0.14, which significantly reduces the computational complexity of the neural network.
-By contrast, the vanilla Transformer architectures in ECCT and FECCT have complexities of $O\left(L\times H\times\left(N^{2}\times d_{k}\times M_{d}+N\times d_{k}^{2}\right)\right)$ and $O\left(L\times H\times\left(N^{2}\times d_{k}+N\times d_{k}^{2}+N^{2}\times M_{d}\right)\right)$, respectively, where $M_{d}$ represents the mask density in these models.
+By contrast, the vanilla Transformer architectures in ECCT and FECCT have complexities of $\mathcal{O}\left(L\times H\times\left(N^{2}\times d_{k}\times M_{d}+N\times d_{k}^{2}\right)\right)$ and $\mathcal{O}\left(L\times H\times\left(N^{2}\times d_{k}+N\times d_{k}^{2}+N^{2}\times M_{d}\right)\right)$, respectively, where $M_{d}$ represents the mask density in these models.
 Table 2: Comparison of trainable parameters, MACs, training time, inference time, and mask density among ECCT, FECCT, and UECCT for various codes, all with identical parameters $L=6$, $H=8$, $d_{k}=64$, $d_{l}=64$, and $d_{f}=2048$

### v4

@@
-Error Correction Code Transformer: From Non-Unified to Unified
+Unified Error Correction Code Transformer with Low Complexity
 Abstract
-Channel coding is vital for reliable data transmission in modern wireless systems, and its significance will increase with the emergence of sixth-generation (6G) networks, which will need to support various error correction codes.
-However, traditional decoders were typically designed as fixed hardware circuits tailored to specific decoding algorithms, leading to inefficiencies and limited flexibility.
-To address these challenges, this paper proposes a unified, code-agnostic Transformer-based decoding architecture capable of handling multiple linear block codes, including Polar, Low-Density Parity-Check (LDPC), and Bose-Chaudhuri-Hocquenghem (BCH), within a single framework.
-To achieve this, standardized units are employed to harmonize parameters across different code types, while the redesigned unified attention module compresses the structural information of various codewords.
-Additionally, a sparse mask, derived from the sparsity of the parity-check matrix, is introduced to enhance the model’s ability to capture inherent constraints between information and parity-check bits, resulting in improved decoding accuracy and robustness.
-Extensive experimental results demonstrate that the proposed unified Transformer-based decoder not only outperforms existing methods but also provides a flexible, efficient, and high-performance solution for next-generation wireless communication systems.
+Channel coding is vital for reliable sixth-generation (6G) data transmission, employing diverse error correction codes for various application scenarios.
+Traditional decoders require dedicated hardware for each code, leading to high hardware costs.
+Recently, artificial intelligence (AI)-driven approaches, such as the error correction code Transformer (ECCT) and its enhanced version, the foundation error correction code Transformer (FECCT), have been proposed to reduce the hardware cost by leveraging the Transformer to decode multiple codes.
+However, their excessively high computational complexity of $\mathcal{O}(N^{2})$ due to the self-attention mechanism in the Transformer limits scalability, where $N$ represents the sequence length.
+To reduce computational complexity, we propose a unified Transformer-based decoder that handles multiple linear block codes within a single framework.
+Specifically, a standardized unit is employed to align code length and code rate across different code types, while a redesigned low-rank unified attention module, with computational complexity of $\mathcal{O}(N)$, is shared across various heads in the Transformer.
+Additionally, a sparse mask, derived from the parity-check matrix’s sparsity, is introduced to enhance the decoder’s ability to capture inherent constraints between information and parity-check bits, improving decoding accuracy and further reducing computational complexity by $86\%$.
+Extensive experimental results demonstrate that the proposed unified Transformer-based decoder outperforms existing methods and provides a high-performance, low-complexity solution for next-generation wireless communication systems.
 1 Introduction
-With the continuous evolution of wireless communication technologies, the imminent arrival of sixth-generation (6G) wireless communications heralds a paradigm shift towards ultra-reliable, low-latency communications and notably enhanced data rates  [1, 2, 3].
-To fulfill the ambitious performance requirements of 6G networks, probabilistic coding schemes are anticipated to play a crucial role.
-In particular, linear block codes such as Polar codes  [4] and Low-Density Parity-Check (LDPC) codes  [5] are expected to maintain their competitive edge established in fifth-generation (5G) wireless communications.
-These codes have also emerged as promising candidates for efficient error correction in 6G  [6].
-Additionally, another algebraic linear block code, known as the Bose-Chaudhuri-Hocquenghem (BCH) code  [7], is often used as an outer code to enhance decoding performance.
+With the continuous evolution of wireless communication technologies, the imminent arrival of sixth-generation (6G) wireless communication will mark a paradigm shift toward ultra-reliable communication [1, 2, 3, 4].
+To meet the stringent performance demands of 6G, channel codes with strong error correction capabilities are expected to play a pivotal role.
+Specifically, low-density parity-check (LDPC) codes [5] and Polar codes [6] have proven effective in fifth-generation (5G) wireless communications.
+Due to their excellent performance, 3GPP has just made the decision to continue to use LDPC and Polar codes for future 6G in October 2025 [7].
+Furthermore, algebraic geometry codes, such as Bose-Chaudhuri-Hocquenghem (BCH) codes [8], are commonly employed as outer codes in concatenated schemes with probabilistic codes to enhance decoding performance.
 1.1 Prior Works
-Looking at 5G and earlier wireless communication systems, linear block code decoders were typically designed as fixed hardware circuits tailored to specific decoding algorithms.
-Looking ahead to 6G, this trend is likely to persist  [8].
-For instance, the Successive Cancellation (SC) decoding algorithm  [4] for Polar codes uses a sequential method that traverses a decision tree to obtain the decoding result.
-This approach simplifies the hardware implementation of Polar codes by pruning binary trees  [9, 10, 11].
-In contrast, the Sum-Product (SP) decoding algorithm  [12] for LDPC codes uses iterative message passing, allowing for parallel data processing.
-This fundamentally differs from the sequential method used for Polar codes in hardware implementation  [13].
-Furthermore, BCH codes typically use algebraic decoding techniques  [14], classified as hard decoding, which differs greatly from soft decoding methods like SC and SP.
-Overall, these algorithms provide strong error correction capabilities.
-However, due to the non-unified nature of the sequential and parallel architectures of Polar and LDPC codes, and the gap between hard and soft decoding in BCH codes, unique hardware circuits must be customized and fine-tuned for each specific linear block code.
-This may lead to significant engineering effort and inefficient use of hardware resources.
-In an effort to reduce hardware resource consumption and establish a quasi-unified decoding framework, three methods have been explored: resource sharing, Ordered-Statistics Decoding (OSD), and Artificial Intelligence (AI)-driven approaches.
-Resource sharing involves time-sharing storage and computing units.
-Ordered-Statistics Decoding leverages the reliability ordering of received symbols to efficiently decode various linear block codes by systematically evaluating the most likely codewords.
-Meanwhile, AI-driven methods offer a convenient way to unfold traditional algorithms like Belief Propagation (BP).
-First, resource-sharing methods, such as those in  [15] and  [16], endeavor to integrate decoders for Polar and LDPC codes into a shared architecture based on traditional decoding algorithms.
-Both approaches aim to conserve hardware resources by temporally multiplexing the computational and storage units shared by the decoder through reconfiguration.
-However, due to the distinct decoding algorithms for each code type, achieving a truly unified architecture with such decoders is generally challenging.
-Second, ordered-statistics decoding methods, initially explored in  [17, 18, 19], provide a near-Maximum Likelihood (ML) decoding approach for linear block codes by exploiting the reliability ordering of received symbols.
-The algorithm operates by sorting the received signal positions based on their reliability, permuting the code’s generator matrix accordingly, and systematically evaluating the most likely codewords through a limited search.
-OSD delivers near-ML performance with significantly less complexity than exhaustive ML decoding, making it well-suited for short to moderate-length codes.
-However, its drawback lies in the computational complexity, which grows exponentially with the search order (e.g., order-$l$).
-For instance, an order-4 OSD may require evaluating thousands of test codewords, potentially leading to increased latency and resource demands.
-Third, in recent years, AI-driven decoders have gained attention as an alternative.
-While model-based AI methods leverage traditional decoder algorithms and rely on neural networks to simulate belief propagation, their performance improvements are often limited by the constraints of conventional models  [20],  [21].
-In contrast, model-free approaches offer a more flexible and scalable solution.
-Unlike model-based methods, which depend on predefined mathematical models, model-free approaches assume no prior knowledge of the encoder or decoder.
-Instead, they rely on large amounts of labeled data to learn the decoding process directly, bypassing the limitations of conventional algorithms  [22].
-For instance, inspired by the success of Transformers in Natural Language Processing (NLP)  [23], a model-free neural decoder called the Error Correction Code Transformer (ECCT) was proposed  [24].
-ECCT employs the Transformer’s self-attention mechanism, where multiple attention heads operate in parallel to process sequential data and capture long-range dependencies, effectively linking input Log-Likelihood Ratios (LLRs) to the decoder’s output.
-The inherent connection between Transformers and decoders lies in their shared ability to translate sequences—whether from one language to another or from LLRs back to the original information bits.
-However, ECCT’s flexibility comes at a cost, requiring frequent parameter adjustments for different code types, lengths, and rates, which increases training complexity and hardware overhead, hindering a unified decoding architecture.
-Building on this,  [25] introduced a Foundation Model for Error Correction Codes (FECCT) to improve ECCT’s generalization to unseen codewords.
-Yet, FECCT currently faces the challenge of high computational complexity in its self-attention mechanism, further complicating its practical implementation.
+In wireless communication systems, multiple coding schemes are supported, such as LDPC and Polar codes in 5G [9], each requiring customized hardware for their decoding algorithms.
+For instance, LDPC codes employ iterative message passing for parallel data processing [10], while Polar codes rely on the successive cancellation (SC) decoding algorithm [6], which follows a sequential approach by traversing a decision tree, often simplified through binary tree pruning [11, 12, 13, 14].
+Both LDPC and Polar codes utilize soft decoding methods, leveraging probabilistic information to achieve robust error correction.
+In contrast, BCH codes, another key coding scheme, employ algebraic decoding techniques classified as hard decoding [15], which rely on mathematical structures to correct errors.The gap between the hard decoding of BCH codes and the soft decoding of LDPC and Polar codes, combined with the distinct parallel and sequential architectures of LDPC and Polar codes, necessitates customized hardware circuits for each code type, resulting in significant hardware resource cost.
+To reduce hardware resource cost, three approaches have been investigated for decoding multiple codecs with a single hardware framework: resource sharing, ordered statistical decoding, and artificial intelligence (AI)-driven methods.
+First, resource-sharing methods, such as those in [16, 17], integrate LDPC and Polar code decoders into a single architecture using traditional decoding algorithms.
+These methods conserve hardware resources by multiplexing shared computational and storage units through reconfiguration.
+However, due to the distinct decoding algorithms for each code type, only partial resource sharing is achieved, limiting hardware resource reduction.
+Second, ordered-statistics decoding methods, as explored in [18, 19, 20], provide near-maximum likelihood (ML) decoding for linear block codes by ordering received symbols by reliability, enabling hardware reuse and reducing resource costs.
+Ordered-statistics decoding sorts signal positions by reliability, permutes the generator matrix of the code in a consistent manner across different linear block codes, and evaluates the most likely codewords through a limited search.
+This uniform approach to generator matrix processing enables hardware reuse across various linear block codes, reducing resource cost compared to code-specific decoding architectures.
+However, the computational complexity of ordered-statistics decoding increases exponentially with code length and search order (e.g., order-$l$).
+For example, an order-4 ordered-statistics decoding may require evaluating thousands of test codewords, increasing decoding latency.
+Third, in recent years, AI-driven decoders have gained attention as an alternative to reduce hardware resource cost.
+For example, inspired by the success of Transformers in natural language processing [21], the error correction code Transformer (ECCT) was introduced to reduce hardware resource and enhance decoding performance for short codes [22].
+The inherent connection between Transformers and decoders lies in their shared capability for sequence-to-sequence transformations, mapping input sequences—e.g., from one language to another in natural language processing or from log-likelihood ratios (LLRs) to information bits in channel decoding—into output sequences.
+ECCT leverages the Transformer’s self-attention mechanism, where multiple attention heads operate in parallel to process sequential data and capture long-range dependencies, effectively modeling the constrained relationships among the LLRs of decoder inputs.
+Building on ECCT, the foundation error correction code Transformer (FECCT) enhances generalization to unseen codewords by integrating the distance matrix, derived from the parity-check matrix $\mathbf{H}$, into the self-attention mechanism [23].
+However, ECCT and FECCT rely on the Transformer’s powerful self-attention mechanism, which has a computational complexity of $\mathcal{O}(N^{2})$, where $N$ represents the sequence length.
+Their high computational complexity makes them difficult to scale to long codes.
 1.2 Contributions
-Thus, to reduce engineering efforts, training costs, and hardware consumption, we propose a unified, code-agnostic Transformer-based decoding architecture capable of handling multiple linear block codes, including Polar, LDPC, and BCH, within a single framework.
-To achieve this, we designed new standardized units and a unified attention module.
-Additionally, a sparse masking mechanism was developed to enhance decoding performance and stability  [^1] Simulation codes will be provided to reproduce the results in this paper: http://oa.ee.tsinghua.edu.cn/dailinglong/publications/publications.html.
+To reduce computational complexity, we propose a unified Transformer-based decoder capable of handling multiple linear block codes, including LDPC, Polar, and BCH, within a single framework.
+To achieve this, we design a standardized unit and a low-rank unified attention module that shares attention scores across various heads in the Transformer.
+Additionally, we develop a sparse masking mechanism to enhance decoding performance and further reduce computational complexity [^1] Simulation codes will be provided to reproduce the results in this paper: http://oa.ee.tsinghua.edu.cn/dailinglong/publications/publications.html..
 Specifically, the main contributions of this paper are summarized as follows.
   1. 1.
-    Unified Decoding Architecture: We introduce a unified, code-agnostic Transformer decoding architecture, capable of seamlessly integrating various linear block codes.
-This architecture features standardized units to align various code lengths and, for the first time, designs a unified attention module that learns to compress the structural information of all codewords, thereby achieving a code-agnostic decoder.
-This architecture obviates the necessity for distinct hardware circuits for different decoding algorithms, reducing engineering efforts and hardware resource inefficiency.
+    We introduce a unified Transformer-based decoder capable of seamlessly integrating various linear block codes.
+This architecture incorporates a standardized unit to align code lengths, a unified attention module that compresses the structural information of all codewords, and a sparse mask leveraging the parity-check matrix’s sparsity to enhance decoding accuracy.
+These components achieve a code-agnostic decoder, eliminate the need for distinct hardware circuits for different decoding algorithms, and reduce computational complexity.
   2. 2.
-    Standardized Units: Standardized units are used to align parameters such as code length and code rate across different code types.
-Specifically, after receiving the channel output, we pad different codewords and their corresponding syndromes with zeros up to the maximum length.
-Through this unit, the Transformer neural decoder can accommodate any codeword within the maximum length, allowing for mixed training in a unified manner, thereby reducing training overhead.
+    A standardized unit aligns parameters, such as code length and code rate, across different code types.
+Specifically, it pads codewords and their corresponding syndromes with zeros to a maximum length after receiving the channel output.
+This enables the Transformer-based decoder to handle any codeword within the maximum length, facilitating unified joint training.
   3. 3.
-    Unified Attention: We redesigned the self-attention module to create a unified attention mechanism that allows different attention heads to share attention scores.
+    We redesigned the self-attention module to create a unified attention mechanism that allows different attention heads to share attention scores.
 This approach contrasts with traditional self-attention mechanisms, where each head independently computes its scores using separate learned weights.
-The unified attention compresses the structural information of different codewords through learning and shares parameters across various heads in the Transformer, resulting in a more cohesive decoding architecture.
+This unified attention compresses structural information of different codewords through learning and shares parameters across various heads in the Transformer.
+It eliminates the need to project inputs into higher dimensions, reducing computational complexity from $\mathcal{O}(N^{2})$ to $\mathcal{O}(N)$, where $N$ represents the sequence length.
+This results in a more efficient and unified decoding architecture.
   4. 4.
-    Performance Enhancement: By introducing a sparse mask, which leverages the sparsity of the parity-check matrix, we have significantly enhanced the model’s ability to capture inherent constraints between information and parity-check bits.
+    By introducing a sparse mask, which leverages the sparsity of the parity-check matrix, we have significantly enhanced the model’s ability to capture inherent constraints between information and parity-check bits.
 This has led to a substantial improvement in decoding performance and the achievement of a state-of-the-art design.
-  5. 5.
-    High Computational Efficiency: The unified attention module in our proposed Transformer-based decoding architecture demonstrates linear computational complexity by eliminating the need to project the input into higher dimensions.
-This results in a significant reduction in complexity from the traditional $\mathcal{O}(N^{2})$ to a more efficient $\mathcal{O}(N)$, where $N$ represents the sequence length.
-Additionally, by employing our sparse mask, we further achieve an $86\%$ reduction in computational complexity.
-This efficiency gain enables faster processing and lower power consumption, making the architecture suitable for real-time applications in next-generation wireless communication systems.
+Additionally, this sparse mask further reduces computational complexity by $86\%$, enabling efficient training for long codes.
 The rest of this paper is organized as follows.
 Section II provides background knowledge, including the vanilla Transformers, channel encoders, as well as pre- and post-processing involved to address the issue of overfitting.
-Section III details our proposed unified code-agnostic Transformer.
+Section III details our proposed unified Transformer-based decoder.
 Section IV presents the results of training and inference.
@@
 The vanilla Transformer, a sequence-to-sequence model, comprises an encoder and decoder, each with $L$ identical blocks.
-Each block integrates multi-head self-attention (MHA), a position-wise feed-forward network (FFN), residual connections  [26] to mitigate gradient issues, and layer normalization  [27] for stability and faster convergence, as shown in Figure 1.
+Each block integrates multi-head self-attention (MHA), a position-wise feed-forward network (FFN), residual connections to mitigate gradient issues [24], and layer normalization for stability and faster convergence [25], as shown in Figure 1.
 Figure 1: Single-layer vanilla Transformer encoder architecture.
 The process begins with embedding a one-dimensional input signal $\textbf{{x}}\in\mathbb{R}^{N}$ into a higher-dimensional space:
-$\textbf{{X}}_{e}=\text{Embedding}(\textbf{{x}})=\text{diag}(\textbf{{x}})\cdot\textbf{{W}}$ (1)
+$\textbf{{X}}_{e}=\text{Embedding}(\textbf{{x}})=\text{diag}(\textbf{{x}})\cdot\textbf{{W}},$ (1)
 where $\text{diag}(\textbf{{x}})\in\mathbb{R}^{N\times N}$ is a diagonal matrix with x on the diagonal, and $\textbf{{W}}\in\mathbb{R}^{N\times d_{k}}$ is a trainable weight matrix, producing $\textbf{{X}}_{e}\in\mathbb{R}^{N\times d_{k}}$ with $d_{k}$-dimensional embeddings for each position.
 This embedding feeds into the MHA mechanism, which correlates sequence elements through:
-$\text{Attn}\left(\textbf{{Q}},\textbf{{K}},\textbf{{V}}\right)=\text{Softmax}\left(\frac{\textbf{{Q}}\cdot\textbf{{K}}^{T}}{\sqrt{d_{k}}}\right)\cdot\textbf{{V}}$ (2)
+$\text{Attn}\left(\textbf{{Q}},\textbf{{K}},\textbf{{V}}\right)=\text{Softmax}\left(\frac{\textbf{{Q}}\cdot\textbf{{K}}^{T}}{\sqrt{d_{k}}}\right)\cdot\textbf{{V}},$ (2)
 extended to multiple heads as:
-$\text{MHA}\left(\textbf{{Q}},\textbf{{K}},\textbf{{V}}\right)=\text{Concat}\left(\text{head}_{1},\cdots,\text{head}_{H}\right)\cdot\textbf{{W}}^{\textit{O}}$ (3)
-where $\text{head}_{i}=\text{Attn}(\textbf{{X}}_{Q}\textbf{{W}}_{i}^{Q},\textbf{{X}}_{K}\textbf{{W}}_{i}^{K},\textbf{{X}}_{V}\textbf{{W}}_{i}^{V})$, with $\textbf{{X}}_{Q},\textbf{{X}}_{K},\textbf{{X}}_{V}\in\mathbb{R}^{N\times d_{k}}$ set to $\textbf{{X}}_{e}$ for self-attention, using trainable weights $\textbf{{W}}_{i}^{Q},\textbf{{W}}_{i}^{K},\textbf{{W}}_{i}^{V}\in\mathbb{R}^{d_{k}\times d_{k}}$ and $\textbf{{W}}^{O}\in\mathbb{R}^{Hd_{k}\times Hd_{k}}$, as depicted in Figure 2.
+$\text{MHA}\left(\textbf{{Q}},\textbf{{K}},\textbf{{V}}\right)=\text{Concat}\left(\text{head}_{1},\cdots,\text{head}_{H}\right)\cdot\textbf{{W}}^{\textit{O}},$ (3)
+where $\text{head}_{i}=\text{Attn}(\textbf{{X}}_{Q}\textbf{{W}}_{i}^{Q},\textbf{{X}}_{K}\textbf{{W}}_{i}^{K},\textbf{{X}}_{V}\textbf{{W}}_{i}^{V})$, with $\textbf{{X}}_{Q},\textbf{{X}}_{K},\\ \textbf{{X}}_{V}\in\mathbb{R}^{N\times d_{k}}$ set to $\textbf{{X}}_{e}$ for self-attention, using trainable weights $\textbf{{W}}_{i}^{Q},\textbf{{W}}_{i}^{K},\textbf{{W}}_{i}^{V}\in\mathbb{R}^{d_{k}\times d_{k}}$ and $\textbf{{W}}^{O}\in\mathbb{R}^{Hd_{k}\times Hd_{k}}$, as depicted in Figure 2.
 Figure 2: Architecture of multi-head self-attention.
 Next, the MHA output is processed by the FFN, applied uniformly across positions:
-$\text{FFN}\left(\textbf{{X}}\right)=\textbf{{W}}_{2}\cdot\text{ReLU}\left(\textbf{{W}}_{1}\cdot\textbf{{X}}+\textbf{{b}}_{1}\right)+\textbf{{b}}_{2}$ (4)
+$\text{FFN}\left(\textbf{{X}}\right)=\textbf{{W}}_{2}\cdot\text{ReLU}\left(\textbf{{W}}_{1}\cdot\textbf{{X}}+\textbf{{b}}_{1}\right)+\textbf{{b}}_{2},$ (4)
 where $\textbf{{W}}_{1}\in\mathbb{R}^{d_{f}\times d_{h}}$, $\textbf{{W}}_{2}\in\mathbb{R}^{d_{h}\times d_{f}}$, $\textbf{{b}}_{1}$, $\textbf{{b}}_{2}$ are trainable parameters, $d_{h}=H\cdot d_{k}$ and $\text{ReLU}(x)=\max(0,x)$ enables complex feature learning, with $d_{f}>d_{h}$ for enhanced performance.
@@
 $\textbf{{X}}_{1}$ $=\text{MHA}\left(\text{LayerNorm}\left(\textbf{{X}}_{e}\right)\right)+\textbf{{X}}_{e}$ (5)
-$\textbf{{X}}_{2}$ $=\text{FFN}\left(\text{LayerNorm}\left(\textbf{{X}}_{1}\right)\right)+\textbf{{X}}_{1}$
+$\textbf{{X}}_{2}$ $=\text{FFN}\left(\text{LayerNorm}\left(\textbf{{X}}_{1}\right)\right)+\textbf{{X}}_{1}.$
 2.2 Forward Error Correction Codes
@@
 The encoder maps a message vector $\textbf{{m}}\in\{0,1\}^{k}$ to a codeword $\textbf{{x}}\in\{0,1\}^{n}$ via $\textbf{{x}}=\textbf{{m}}\cdot\textbf{G}$, where $\textbf{H}\cdot\textbf{{x}}=\textbf{0}$.
-This codeword is modulated using Binary Phase Shift Keying (BPSK), defined as $\text{BPSK}(x)=1-2x$, yielding $\textbf{{x}}_{s}$, which is transmitted over Binary-Input Discrete Memoryless Channels (B-DMCs, e.g., AWGN), producing $\textbf{{y}}=\textbf{{x}}_{s}+\textbf{{n}}$, with $\textbf{{n}}\sim\mathcal{N}(0,\sigma^{2})$.
+This codeword is modulated using binary phase shift keying (BPSK), defined as $\text{BPSK}(x)=1-2x$, yielding $\textbf{{x}}_{s}$, which is transmitted over binary-input discrete memoryless channels (B-DMCs, e.g., additive white Gaussian noise (AWGN) channels), producing $\textbf{{y}}=\textbf{{x}}_{s}+\textbf{{n}}$, with $\textbf{{n}}\sim\mathcal{N}(0,\sigma^{2})$.
 Figure 3: Architecture of the communication system.
@@
 Our work focuses on designing and training a parameterized decoding function $f_{\theta}$.
-To develop a model-free Transformer-based decoder, we leverage pre- and post-processing techniques from  [28], preserving BER and Mean Squared Error (MSE) performance (see Figure 4).
+To develop a data-driven Transformer-based decoder, we leverage pre- and post-processing techniques from [26], preserving BER and mean squared error (MSE) performance (see Figure 4).
 Training with an all-zero codeword mitigates overfitting due to exponential codeword space growth.
 And the B-DMCs is modeled by (6) which is an equivalent statistical mode that differs from the true physical one [27].
-$\textbf{{y}}=\textbf{{x}}_{s}\cdot\textbf{{z}}$ (6)
+$\textbf{{y}}=\textbf{{x}}_{s}\cdot\textbf{{z}},$ (6)
 where z is a random multiplicative noise independent of $\textbf{{x}}_{s}$}.
 In the pre-processing step, the absolute values of received signal and the syndromes are concatenated as the input to the neural network, which can be written as:
-$\widetilde{\textbf{{y}}}=\left[\lvert\textbf{{y}}\rvert,\text{s}\left(\textbf{{y}}\right)\right]$ (7)
+$\widetilde{\textbf{{y}}}=\left[\lvert\textbf{{y}}\rvert,\text{s}\left(\textbf{{y}}\right)\right],$ (7)
 where, $\left[\cdot,\cdot\right]$ denotes the concatenation of vectors, $\lvert\textbf{{y}}\rvert$ signifies the absolute value of y and $\text{s}(\textbf{{y}})=\textbf{H}\cdot\text{bin}(\text{sign}(\textbf{{y}}))\in\{0,1\}^{n-k}$ is the syndrome from hard-decoded y.
@@
 3 Proposed Unified Error Correction Code Transformer
-In this section, we present the complete architecture and training process of the proposed Unified Error Correction Code Transformer (UECCT).
+In this section, we present the detailed architecture and training process of the proposed unified error correction code Transformer (UECCT).
 3.1 Shared Unified Attention
@@
 This approach, while effective for capturing long-range dependencies, is computationally intensive and overlooks the intrinsic relationships across different samples, particularly in channel decoding tasks involving diverse linear block codes.
-Moreover, prior studies [30] suggest that the input to the softmax function in self-attention is probabilistically sparse, with only a few weights dominating the attention distribution.
+Moreover, prior studies suggest that the input to the softmax function in self-attention is probabilistically sparse, with only a few weights dominating the attention distribution [28].
 This sparsity implies that a low-rank approximation could effectively capture the dominant correlations, reducing redundancy without compromising performance.
 Figure 5: Rank distribution of the $\textbf{{Q}}\cdot\textbf{{K}}^{T}$ matrix in the ECCT model across 6 encoder layers for various code types.
-To investigate this hypothesis, we trained an ECCT model [24] with parameters $L=6$ (number of Transformer encoder layers), $H=8$, $d_{k}=\text{max}(2n-k)=182$, and $d_{f}=2912$ on a diverse set of linear block codes, including Polar(32,16), Polar(64,32), Polar(128,86), LDPC(49,24), LDPC(121,60), LDPC(121,80), BCH(31,16), BCH(63,36), and BCH(127,120).
+To investigate this hypothesis, we trained an ECCT model [22] with parameters $L=6$ (number of Transformer encoder layers), $H=8$, $d_{k}=\text{max}(2n-k)=182$, and $d_{f}=2912$ on a diverse set of linear block codes, including LDPC(49,24), LDPC(121,60), LDPC(121,80), Polar(32,16), Polar(64,32), Polar(128,86), BCH(31,16), BCH(63,36), and BCH(127,120).
 We analyzed the rank distribution of the $\textbf{{Q}}\cdot\textbf{{K}}^{T}$ matrix across 1000 batches for the six-layer Transformer encoder.
@@
 However, we hypothesize that the attention distributions across heads may exhibit high similarity, particularly in channel decoding, where the sparse structure of the parity-check matrix $\mathbf{H}$ constrains bit interactions.
-To quantify this, we employed the Jensen-Shannon Divergence (JSD) [31], a symmetric measure of similarity between probability distributions, defined as:
-$\text{JSD}(\mathbf{A}_{i}[k]\|\mathbf{A}_{j}[k])=\frac{1}{2}\left[D_{KL}(\mathbf{A}_{i}[k]\|\mathbf{M}[k])+D_{KL}(\mathbf{A}_{j}[k]\|\mathbf{M}[k])\right]$ (9)
+To measure attention head similarity, we employed the Jensen-Shannon Divergence (JSD) [29], a symmetric measure of similarity between probability distributions, defined as:
+$\text{JSD}(\mathbf{A}_{i}[k]\|\mathbf{A}_{j}[k])=\frac{1}{2}\left[D_{KL}(\mathbf{A}_{i}[k]\|\mathbf{M}[k])+D_{KL}(\mathbf{A}_{j}[k]\|\mathbf{M}[k])\right],$ (9)
 where $k$ is the row index, $\mathbf{A}_{i}$ and $\mathbf{A}_{j}$ $\in\mathbb{R}^{N\times N}$are the softmax outputs of attention heads $i$ and $j$, $\mathbf{M}[k]=\frac{1}{2}(\mathbf{A}_{i}[k]+\mathbf{A}_{j}[k])$ is the average distribution, and $D_{KL}(\mathbf{P}\|\mathbf{Q})=\sum_{l=1}^{N}\mathbf{P}[l]\log\frac{\mathbf{P}[l]}{\mathbf{Q}[l]}$ is the Kullback-Leibler divergence [30].
@@
 The UAM is defined as:
-$\text{UniAttn}\left({{\textbf{{A}}_{l}},{\textbf{{V}}_{l}}}\right)=\text{Softmax}\left({{\textbf{{A}}_{l}}}\right)\cdot\left({{\textbf{{V}}_{l}^{T}}\textbf{{X}}}\right)$ (10)
+$\text{UniAttn}\left({{\textbf{{A}}_{l}},{\textbf{{V}}_{l}}}\right)=\text{Softmax}\left({{\textbf{{A}}_{l}}}\right)\cdot\left({{\textbf{{V}}_{l}^{T}}\textbf{{X}}}\right),$ (10)
 where $\textbf{{X}}\in\mathbb{R}^{N\times d_{k}}$ is the input to the attention layer. In contrast, the conventional self-attention is expressed as:
-$\text{Attn}\left(\textbf{{W}}^{\textit{Q}},\textbf{{W}}^{\textit{K}},\textbf{{W}}^{\textit{V}}\right)=\text{Softmax}\left(\frac{\left(\textbf{{XW}}^{\textit{Q}}\right)\cdot{\left(\textbf{{XW}}^{\textit{K}}\right)}^{T}}{\sqrt{d_{k}}}\right)\cdot\left(\textbf{{XW}}^{\textit{V}}\right)$ (11)
+$\text{Attn}\left(\textbf{{W}}^{\textit{Q}},\textbf{{W}}^{\textit{K}},\textbf{{W}}^{\textit{V}}\right)=\text{Softmax}\left(\frac{\left(\textbf{{XW}}^{\textit{Q}}\right)\cdot{\left(\textbf{{XW}}^{\textit{K}}\right)}^{T}}{\sqrt{d_{k}}}\right)\cdot\left(\textbf{{XW}}^{\textit{V}}\right).$ (11)
 By eliminating the projections $\textbf{{W}}^{\textit{Q}}$ and $\textbf{{W}}^{\textit{K}}$, the UAM reduces the computational complexity from $\mathcal{O}(N^{2}\cdot d_{k}+N\cdot d_{k}^{2})$ to $\mathcal{O}(N\cdot d_{l}\cdot d_{k})$, where $d_{l}\ll N$.
-The shared attention matrix $\mathbf{A}_{l}$ is applied across all heads, leveraging the observed head similarity to foster a cohesive decoding architecture.
+The shared attention matrix $\mathbf{A}_{l}$ is applied across all heads, leveraging the observed head similarity to promote a unified decoding architecture.
 This design not only reduces the parameter count but also enables the model to learn shared features across diverse code types, facilitating a code-agnostic framework.
@@
 The UAM is formulated as:
-$\text{UniAttn}\left(\boldsymbol{A}_{l},\boldsymbol{V}_{l}\right)=\text{Softmax}\left(\boldsymbol{A}_{l}+\text{M}(\textbf{H})\right)\cdot\left(\boldsymbol{V}_{l}^{T}\boldsymbol{X}\right)$ (12)
+$\text{UniAttn}\left(\boldsymbol{A}_{l},\boldsymbol{V}_{l}\right)=\text{Softmax}\left(\boldsymbol{A}_{l}+\text{M}(\textbf{H})\right)\cdot\left(\boldsymbol{V}_{l}^{T}\boldsymbol{X}\right).$ (12)
 The design of the mask function $\text{M}(\textbf{H})$ aims to enable the neural network to capture the relationship between the received signal $\textbf{{y}}\in\mathbb{R}^{n}$ and its syndrome $\text{s}(\textbf{{y}})=\textbf{H}\cdot\overline{\textbf{{y}}}\in\{0,1\}^{n-k}$, where $\overline{\textbf{{y}}}=\text{bin}(\text{sign}(\textbf{{y}}))$ is the hard-decoded binary vector derived from y.
@@
 To embed this relationship within the attention mechanism, we define an extended parity-check matrix $\overline{\textbf{H}}\in\mathbb{R}^{(2n-k)\times(n-k)}$ satisfying:
-$\left[\overline{\textbf{{y}}},\text{s}\left(\textbf{{y}}\right)\right]\cdot\overline{\textbf{H}}=\textbf{0}$ (13)
+$\left[\overline{\textbf{{y}}},\text{s}\left(\textbf{{y}}\right)\right]\cdot\overline{\textbf{H}}=\textbf{0}.$ (13)
 Inspired by this, we explore a mask structure leveraging the parity-check constraints.
 Let $\textbf{{A}}_{l}$ be a matrix in $\mathbb{R}^{N\times d_{l}}$, where $N$ is defined as 2nk2n-k and $d_{l}$ as nkn-k.
-The theoretical foundation for setting $d_{l}\approx n-k$ stems from the maximum rank of the parity-check matrix $\textbf{H}\in\mathbb{F}_{2}^{(n-k)\times n}$, which is nkn-k.
+The theoretical foundation for setting $d_{l}=n-k$ stems from the maximum rank of the parity-check matrix $\textbf{H}\in\mathbb{F}_{2}^{(n-k)\times n}$, which is nkn-k.
 We then construct the extended parity-check matrix $\overline{\textbf{H}}$ as follows:
-$\overline{\textbf{H}}=\begin{bmatrix}{\textbf{H}}^{T}\\ {\textbf{{I}}}_{n-k}\end{bmatrix}$ (14)
+$\overline{\textbf{H}}=\begin{bmatrix}{\textbf{H}}^{T}\\ {\textbf{{I}}}_{n-k}\end{bmatrix}$, (14)
 where ${\textbf{{I}}}_{n-k}\in\mathbb{R}^{(n-k)\times(n-k)}$ is an identity matrix with ones on the diagonal and zeros elsewhere.
@@
 We naturally proceed to define the density of $\overline{\textbf{H}}$ as:
-$H_{d}=\frac{\text{sum}(\overline{\textbf{H}})}{(2n-k)\times(n-k)}$ (15)
+$H_{d}=\frac{\text{sum}(\overline{\textbf{H}})}{(2n-k)\times(n-k)}.$ (15)
 Figure 8: We present the corresponding matrices $\textbf{H}^{T}$, $\overline{\textbf{H}}$ and the Tanner graph for the Hamming (7,4) code.
@@
 Here, matrix H is defined as:
-$\textbf{H}=\begin{bmatrix}1&1&1&0&1&0&0\\ 1&0&1&1&0&1&0\\ 0&1&1&1&0&0&1\end{bmatrix}$ (16)
+$\textbf{H}=\begin{bmatrix}1&1&1&0&1&0&0\\ 1&0&1&1&0&1&0\\ 0&1&1&1&0&0&1\end{bmatrix}$. (16)
 Algorithm 2 Pseudocode for Sparse Mask Construction
@@
 This illustration demonstrates that the application of the mask results in sparser attention, a phenomenon particularly pronounced in low-density parity-check codes.
-Most importantly, the complexity of the UAM is now reduced to the density of the code $\mathcal{O}(N\times d_{l}\times d_{k}\times H_{d})$.
+Most importantly, the computational complexity of the UAM is now reduced to the density of the code $\mathcal{O}(N\times d_{l}\times d_{k}\times H_{d})$.
 Our experimental results indicate that the use of this mask has successfully reduced the computational complexity of the attention mechanism by an average of 86%.
 Furthermore, we provide a summary of the Python-style sparse mask construction within Algorithm 2.
-It should be noted that, in Application-Specific Integrated Circuit (ASIC) implementations, computations involving the masked parts can be directly skipped to enhance efficiency.
+It should be noted that, in application-specific integrated circuit (ASIC) implementations, computations involving the masked parts can be directly skipped to enhance efficiency [31].
 3.3 Architecture and Training Methodology
@@
 The primary innovation in our model involves substituting the vanilla attention mechanism with a multi-head shared unified-attention module.
-This modification not only decreases computational complexity but also empowers the model to discern distinctive features among diverse codewords, thereby facilitating the establishment of a unified decoding architecture.
+This modification not only decreases computational complexity but also empowers the model to identify distinctive features among diverse codewords, thereby facilitating the establishment of a unified decoding architecture.
 We employ fine-tuning techniques to enhance the model’s generalization to unseen codewords.
@@
 Therefore, the loss calculation for a received individual codeword y is given by:
-$\text{loss}=-\sum_{i=1}^{n}{\tilde{z}}_{i}\text{log}(f_{\theta}(y))+(1-{\tilde{z}}_{i})\text{log}(1-f_{\theta}(y))$ (17)
+$\text{loss}=-\sum_{i=1}^{n}{\tilde{z}}_{i}\text{log}(f_{\theta}(y))+(1-{\tilde{z}}_{i})\text{log}(1-f_{\theta}(y)).$ (17)
 To accommodate codewords of varying lengths within a unified architecture, we have designed a standardization unit that performs padding on $\lvert\textbf{{y}}\rvert$ and $s(\textbf{{y}})$ as described in Equation (7).
@@
 Assuming that the maximum length of the codeword is $N_{\text{max}}$ and the maximum length of the syndrome is $S_{\text{max}}$, the dataset is standardized according to:
-$\widetilde{\textbf{{y}}}=\left[~\left[\lvert\textbf{{y}}\rvert,~\textbf{0}_{c}\right],~\left[s(\textbf{{y}}),~\textbf{0}_{s}\right]~\right]$ (18)
+$\widetilde{\textbf{{y}}}=\left[~\left[\lvert\textbf{{y}}\rvert,~\textbf{0}_{c}\right],~\left[s(\textbf{{y}}),~\textbf{0}_{s}\right]~\right],$ (18)
 where $\textbf{0}_{c}$ and $\textbf{0}_{s}$ represent zero vectors with sizes $N_{\text{max}}-\text{size}(\textbf{{y}})$ and $S_{\text{max}}-\text{size}(s(\textbf{{y}}))$, respectively.
-Through this unit, the Transformer neural decoder can accommodate any codeword within the maximum length, allowing for mixed training in a unified manner.
-In ASIC implementations, computations for zero-padded parts can be skipped to improve efficiency, similar to the sparse-masked attention.
+Through this unit, the Transformer neural decoder can accommodate any codeword within the maximum length, allowing for joint training in a unified manner.
+In ASIC implementations, computations for zero-padded parts can be skipped to improve efficiency, similar to the sparse-masked attention [31].
 In the training phase, we conducted 1000 epochs, each consisting of 1000 minibatches, which in turn contained 512 samples each.
-Samples for each minibatch were randomly selected from a pool of systematic Polar [33], LDPC, and BCH codewords.
+Samples for each minibatch were randomly selected from a pool of LDPC, Polar, and BCH codewords.
 The learning rate was initialized at $10^{-3}$, and a cosine decay scheduler was applied without warmup [32], gradually reducing it to $10^{-6}$ by the end of the training process.
@@
 4 Experiments
-We conducted experiments on linear block codes, specifically Polar, LDPC, and BCH, to assess the effectiveness of the proposed architecture.
+We conducted experiments on linear block codes, specifically LDPC, Polar, and BCH, to assess the effectiveness of the proposed architecture.
 The parity-check matrices utilized in the paper were sourced from [34].
-We benchmark our approach against the most recent state-of-the-art performance from  [24], which trains each codeword parameter independently, and the hybrid-trained FECCT  [25].
-Additionally, we compare it to legacy BP-based decoders, including unlearned BP  [37], fully supervised Hyper BP  [38], and Autoregressive BP (AR BP)  [39].
+We benchmark our approach against the most recent state-of-the-art methods from [22], which trains each codeword parameter independently, and the jointly trained FECCT [23].
+Additionally, we compare it to legacy belief propagation (BP)-based decoders, including unlearned BP [35], fully supervised Hyper BP [36], and Autoregressive BP (AR BP) [37].
 The results are presented in the form of bit error rate (BER) and block error rate (BLER) at various normalized SNR (i.e.
-$E_{b}/N_{0}$) values, with at least $10^{5}$ random codewords tested per code type.
+$\mathrm{E_{b}/N_{0}}$) values, with at least $10^{5}$ random codewords tested per SNR.
 We define aggregate BER (ABER) and aggregate BLER (ABLER) as the overall error rates spanning all evaluated codewords.
-In addition, to enable direct comparison with ECCT and FECCT, whose simulation results were directly obtained from their respective papers  [24, 25], we fine-tuned the pre-trained UECCT models on Consultative Committee for Space Data Systems (CCSDS)  [40] and MacKay  [41] codewords leveraging the hybrid-trained Polar, LDPC, and BCH models.
+In addition, to enable direct comparison with ECCT and FECCT, whose simulation results were sourced from their respective papers [22, 23], we fine-tuned the pre-trained UECCT models on consultative committee for space data systems (CCSDS) [38] and MacKay [39] codewords using jointly trained LDPC, Polar, and BCH models.
 Given GPU memory constraints and the fact that the performance of long codes has been thoroughly explored [40], while the error correction of short codes remains far from the Shannon limit [41], this work primarily focuses on training and validating medium-to-short codes.
@@
 We present the performance of our framework with hyperparameters are set to $L=6$, $H=8$, $d_{k}=64$, $d_{l}=64$, and $d_{f}=4\cdot H\cdot d_{k}$.
-The value of $d_{l}=64$ corresponds to the maximum syndrome length among the trained codewords.
+The value of $d_{l}=64$ corresponds to the maximum syndrome length among the jointly trained codewords.
+Table 1: Negative natural logarithm of BER for ECCT, FECCT, and UECCT across Polar, LDPC, BCH, CCSDS, and MacKay codes at $\mathrm{E_{b}/N_{0}}$ = 4, 5, 6 dB, with best results in bold (higher is better).
 Figure 10: BER comparison of ECCT, FECCT, and proposed UECCT for Polar codes.
@@
 Figure 12: BER comparison of ECCT, FECCT, and proposed UECCT for BCH codes.
-Table 1: A comparison of the negative natural logarithm of the BER
-Table 1 presents the simulation results, specifically showing the negative natural logarithm of the BER at $E_{b}/N_{0}$ values of 4, 5, and 6 dB.
+Table 1 presents the simulation results, specifically showing the negative natural logarithm of the BER at $\mathrm{E_{b}/N_{0}}$ values of 4, 5, and 6 dB.
 It is noted that higher values in this context indicate better performance, with the best results for each SNR highlighted in bold.
 Additional BER plots for Polar, LDPC and BCH codes are provided in Figure 10, 11 and Figure 12, respectively.
-Simulation results demonstrate that our proposed UECCT outperforms both the individually trained ECCT  [24] and the hybrid-trained FECCT  [25] across the majority of codewords, validating the effectiveness of the proposed architecture.
-Notably, UECCT even surpasses FECCT’s performance with an embedding length of 128 when operating at a more compact embedding length of 64, further highlighting the efficiency and power of our design.
-The superior performance stems from its ability to leverage a shared attention mechanism across heads, as detailed in Section 3, combined with sparsity-aware masking that reduces computational complexity while preserving decoding accuracy.
+Simulation results demonstrate that our proposed UECCT outperforms both the individually trained ECCT [22] and the jointly trained FECCT [23] across the majority of codewords, validating the effectiveness of the proposed architecture.
+Notably, UECCT even surpasses FECCT’s performance with an embedding length of 128 when operating at a more compact embedding length of 64, further highlighting the computational efficiency of our design.
+The superior performance arises from its ability to utilize a shared attention mechanism across heads, combined with sparsity-aware masking that reduces computational complexity while maintaining decoding accuracy.
+Figure 13: BER comparison between the NMS algorithm and the proposed UECCT for medium-to-long LDPC codes.
+To further validate the performance and scalability of the proposed method on medium-to-long codes, we conducted joint training using LDPC(576, 288), LDPC(768, 512), and LDPC(960, 800) codewords with hyperparameters set to $L=12$, $H=8$, $d_{k}=16$, $d_{l}=288$, and $d_{f}=512$.
+The neural network training and inference were performed under an AWGN channel with BPSK modulation.
+During training, the SNR was randomly selected between 2 dB and 3 dB, with a batch size of 192, while other training methods were consistent with those described in Section 3.3.
+During inference, $10^{5}$ code blocks were tested per SNR.
+For comparison with traditional decoding algorithms, we utilized the normalized min-sum (NMS) algorithm for LDPC codes [42].
+The NMS algorithm was configured with a normalization factor of 0.875 and a maximum of 15 iterations.
+Additionally, an early stopping mechanism was implemented to halt decoding once the parity-check matrix $\mathbf{H}$ is satisfied.
+Figure 13 presents the BER performance comparison between the NMS algorithm and the proposed UECCT.
+The simulation results demonstrate that the proposed method achieves performance comparable to the traditional NMS decoding algorithm.
+This indicates that our decoding architecture effectively scales to medium-to-long codes.
 5 Analysis
-We investigate the impact of Transformer encoder layers on performance, focusing on the embedding dimension $d_{k}$ and the effectiveness of sparse masked attention.
-Additionally, we compare the computational complexity of our approach with existing methods.
+Using the parameters in Table 1, we examine the effect of Transformer encoder layers on performance, focusing on the embedding dimension $d_{k}$ and sparse masked attention efficiency.
+We also compare our approach’s computational complexity with that of existing methods.
 5.1 Impact of Encoder Layers and Embedding Dimensions
@@
 Deeper architectures better capture intricate data patterns and dependencies, crucial for error correction.
-Additional layers enable multi-level abstraction and refinement, improving error interpretation and correction, especially in complex wireless communication systems where simpler models falter.
+Additional layers enable multi-level abstraction and refinement, improving error estimation and correction, especially in complex wireless communication systems where simpler models falter.
 These findings highlight the primacy of model depth in Transformer-based error correction. Prioritizing $L$ over $d_{k}$ can optimize model design, enhancing practical performance.
@@
 By contrast, the vanilla Transformer architectures in ECCT and FECCT have complexities of $\mathcal{O}\left(L\times H\times\left(N^{2}\times d_{k}\times M_{d}+N\times d_{k}^{2}\right)\right)$ and $\mathcal{O}\left(L\times H\times\left(N^{2}\times d_{k}+N\times d_{k}^{2}+N^{2}\times M_{d}\right)\right)$, respectively, where $M_{d}$ represents the mask density in these models.
-Table 2: Comparison of trainable parameters, MACs, training time, inference time, and mask density among ECCT, FECCT, and UECCT for various codes, all with identical parameters $L=6$, $H=8$, $d_{k}=64$, $d_{l}=64$, and $d_{f}=2048$
+Table 2: Comparison of trainable parameters, MACs, training time, inference time, and mask density among ECCT, FECCT, and UECCT for various codes, all with identical parameters $L=6$, $H=8$, $d_{k}=64$, $d_{l}=64$, and $d_{f}=2048$.
 Table 2 compares the trainable parameters (TPs), multiply-accumulate operations (MACs), training time, inference time, and mask density for ECCT, FECCT, and UECCT.
@@
 This inefficiency could limit its potential for deployment.
-To address these issues, future implementations could leverage techniques such as parameter sharing, low-rank factorization, and quantization, aiming to reduce complexity and enhance efficiency  [45, 46].
+To address these issues, future implementations could leverage techniques such as parameter sharing, low-rank factorization, and quantization, aiming to reduce complexity and enhance efficiency [44, 45, 46].
 6 Conclusions
-We introduced a unified, code-agnostic Transformer decoding architecture, designed to integrate various linear block codes seamlessly and offer superior decoding performance.
-Our approach leverages the self-attention mechanism inherent in Transformer models, enabling effective handling of sequential data and capturing long-range dependencies critical for decoding tasks.
-Through rigorous analysis and extensive experimental validation, our results indicate that the unified Transformer decoding architecture can significantly advance the state-of-the-art in decoding performance, potentially setting new benchmarks for future research and practical applications.
-In conclusion, the unified Transformer decoding architecture presented in this paper signifies a significant step forward in the field of error correction coding.
-It paves the way for the development of more adaptable, efficient, and robust decoding solutions, poised to meet the stringent requirements of 6G and beyond.
-As the field progresses, we anticipate that the continued integration of AI technologies will unlock new possibilities, driving innovation in wireless communication systems.
-Future work can aim to further reduce hardware implementation complexity and enhance computational efficiency using techniques such as quantization and compression of neural network models.
+We proposed a unified Transformer-based decoder that seamlessly integrates multiple linear block codes within a single framework.
+By introducing standardized unit, a unified attention module, and a sparse mask leveraging the parity-check matrix’s sparsity, we achieved enhanced decoding accuracy and reduced computational complexity from $\mathcal{O}(N^{2})$ to $\mathcal{O}(N)$.
+This work delivers a high-performance, low-complexity solution, addressing hardware overhead and scalability challenges in next-generation wireless systems like 6G.
+Future work will focus on improving memory efficiency and power consumption through techniques such as pruning and quantization to enhance deployability in resource-constrained environments.